### v1

# Personalized Item Embeddings in Federated Multimodal Recommendation

Abstract.
Federated recommendation systems play a crucial role in protecting user privacy. However, existing methods primarily rely on ID-based item embeddings, overlooking the rich multimodal information of items. To address this limitation, we propose a novel Federated Multimodal Recommendation System called FedMR. FedMR leverages a foundation model on the server side to encode multimodal data, such as images and text, associated with items. To tackle the challenge of data heterogeneity caused by varying user preferences, FedMR introduces a Mixing Feature Fusion Module on the client. This module dynamically adjusts the weights of different fusion strategies based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences. FedMR is compatible with existing ID-based federated recommendation systems, improving their performances without modifying the original framework. Our experiments on four real-world multimodal recommendation datasets demonstrate the effectiveness of FedMR. Our code is available at https://anonymous.4open.science/r/FedMR.
Federated Recommendation, Multimodal Learning, Personalization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
1. Introduction
Under privacy protection regulations, such as GDPR (Voigt and Von dem Bussche, 2017), that emphasize user data privacy, Federated Recommendation Systems (FedRec) have become an essential technology in modern life (Sun et al., 2024). However, in these systems, since each client typically only has access to a single user’s interaction history, data sparsity is a common issue that degrades recommendation performance (Li and Long, 2024). Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity. Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item feature representations can significantly improve recommendation results. Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data. Thus, generating personalized item embeddings using multimodal data while preserving user privacy remains a significant challenge.
As a potential solution, Multimodal Recommendation Systems (MMRec) integrate data from different modalities, significantly enhancing recommendation performance (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a). Unlike single-modal systems, MMRec can capture hidden relationships and complementary information between various modalities, leading to a better understanding of user preferences and item characteristics (Zhou et al., 2023b). For example, MMRec methods (Wei et al., 2019; Wang et al., 2021; Tao et al., 2022) combine video frames, audio tracks, and item descriptions when users interact with online platforms, providing a more accurate representation of user preferences. To effectively handle multimodal data, MMRec often uses Foundation Models (FM) pretrained on large-scale datasets, which possess powerful feature extraction and representation capabilities (Bommasani et al., 2021). By applying FMs, MMRec can efficiently extract and integrate information from multiple modalities, capturing commonalities and complementary relationships between them (Radford et al., 2021a; Cui et al., 2022; Geng et al., 2022, 2023; Wu et al., 2024). However, these centralized MMRec methods rely on servers to aggregate user data, digital activities, and preferences to train accurate models (Liu et al., 2024), which raises privacy concerns (Chai et al., 2020; Li et al., 2023a).
To address these challenges, Federated Multimodal Recommendation Systems (FedMMRec) have emerged, which integrates MMRec and introduces the capability of processing multimodal information into FedRec. By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024). FedMMRec is still in the early stages of research, and related studies are limited. FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering. AMMFRS (Feng et al., 2024) introduces an attention mechanism on the client side to perform weighted fusion of multimodal data, effectively distinguishing and utilizing the contributions of different modalities. These methods not only protect user privacy but also effectively integrate multimodal features to improve recommendation accuracy and personalization. However, they rely on fixed multimodal fusion strategies, neglecting the personalized representation of item features on clients, which limits their ability to adapt to diverse user needs. Although AMMFRS uses an attention mechanism to distinguish the importance of different modalities, it lacks adaptability to the diversity of user data, resulting in limited flexibility and accuracy of fusion strategies (Niu et al., 2021).
Main Contribution. To integrate multimodal item information into FedRec while ensuring user privacy, we propose a Federated Multimodal Recommendation system (FedMR) based on foundation models. FedMR uses a foundation model on the server side to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items. On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts fusion strategies based on user interaction history, combining multimodal and ID embeddings to generate personalized item embeddings. This approach provides more personalized recommendations. Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients.
The main contributions of this work are summarized as follows:
  • We propose FedMR, a federated multimodal recommendation system using foundation models, which efficiently leverages multimodal data while preserving user privacy;
  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item embeddings, enhancing personalization and flexibility in recommendations;
  • FedMR is compatible with existing ID-based federated recommendation methods, improving recommendation performance without modifying the existing framework;
  • Extensive experiments are conducted on four real-world multimodal recommendation datasets to validate the effectiveness of our proposed method FedMR;
  • We provide the source code to support the reproducibility.
2. Related Work
2.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely heavily on ID embeddings to model items, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item embeddings, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
2.2. Multimodal Recommendation Systems
Unlike traditional Recommendation Systems (RecSys) that rely primarily on user-item interaction data (such as clicks or ratings) to capture user behavior (Yu et al., 2023), Multimodal Recommendation systems (MMRec) aim to improve the accuracy of recommendations and enhance user experience by utilizing multiple modalities of data , such as text, images, audio, and video (Zhou et al., 2023b). As unique identifiers for items, item IDs not only contain semantic information about the items but also capture the characteristics of neighboring items (Yuan et al., 2023b; Liu et al., 2023). Therefore, in MMRec, item IDs are often used as auxiliary information and combined with multimodal data to enhance item representations(Wang et al., 2024; Zhang et al., 2024e). This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving the quality of recommendations. A typical MMRec training process generally involves feature extraction and modality fusion (He and McAuley, 2016; Wei et al., 2019; Wang et al., 2021; Tao et al., 2022). Feature extraction (Guo et al., 2019) aims to obtain representative features from multiple modalities to better capture the attributes and preferences of users and items. Modality fusion (Gao et al., 2020) focuses on effectively integrating data from different modalities to enhance the overall performance of the MMRec and the relevance of its results. The recommendation models (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023) combine multimodal data with user-item interaction data to predict user preferences and provide personalized recommendations.
2.3. Foundation Models
Foundation Models (FM) are often pre-trained on large-scale unlabeled datasets for general purposes, that can be adapted to a variety of downstream tasks through appropriate fine-tuning, such as task-specific training (Bommasani et al., 2021). This makes them highly generalizable and adaptable across different applications. They are then adapted to specific recommendation tasks through methods such as representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), model fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), and prompt tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024). Some studies have explored the use of FM to enhance the content-awareness of MMRec, improving content understanding and generation by leveraging alignment and generation techniques across multiple modalities. CLIP (Radford et al., 2021a; Wu et al., 2024) uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain multimodal representations, capturing the matching relationships between users and items. M6-Rec (Cui et al., 2022) builds a recommendation system based on large-scale multimodal pre-trained models specifically designed for the e-commerce domain. P5 (Geng et al., 2022) transforms the recommendation task into a natural language processing task, generating personalized recommendations through prompts. VIP5 (Geng et al., 2023) leverages pre-trained multimodal models to capture complementary information from visual and textual data, providing more personalized and relevant recommendations to users. FedPA (Zhang et al., 2024c) and TransFR (Zhang et al., 2024b) integrate FM with FedRec. FedPA introduces a personalized adapter to fine-tune the FM, allowing each client to learn individual user preferences in a lightweight manner. In contrast, TransFR leverages FMs to generate universal textual representations, replacing traditional ID-based item representations and enhancing the system’s adaptability to new domains and its ability to handle cold-start issues.
3. Problem Formulation
Let $\mathcal{U}=\{u\}$ and =$\mathcal{I}=\{i\}$ be the set of users and items correspondingly, we assume that the users’ interaction history data is implicit feedback, represented by $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n}]^{T}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|},$ where $r_{u,i}=1$ if user $u$ has interacted with item $i$, and $r_{u,i}=0$ otherwise. For user $u$, we use $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$ to denote their interaction history and $\mathcal{I}_{u}$ to represent the set of items they have interacted with. Moreover, we define $\mathcal{O}=\{(u,i)|\text{user }u\text{ has interacted with item }i\}$ to index the items each user has interacted with. Each item in $\mathcal{I}$ has a set of modality features $\mathcal{M}$. The $m$-th modality feature is represented as $\mathbf{E}_{m}=[\mathbf{e}_{1,m},\mathbf{e}_{2,m},\dots,\mathbf{e}_{|\mathcal{I}|,m}]$. To capture user diversity, FedMR also introduces a set of modality fusion strategies $\mathcal{G}=\{g_{1},g_{2},\cdots,g_{|\mathcal{G}|}\}$. In this work, we focus on the visual and textual modalities, denoted as =$\mathcal{M}=\{v,c\}$, respectively. It is important to note that FedMR is not limited to these two modalities and the existing fusion strategies, it can include additional modalities or modify the fusion strategies as needed.
3.1. Objective of FedRec
In classic FedRec with implicit feedback, each client only has the user’s historical interaction data $\mathbf{r}_{u}$. The goal is to train a personalized recommendation model for each user while protecting user privacy, providing personalized recommendations by predicting possible scores $\hat{\mathbf{r}}_{u}$ for items the user has not yet interacted with, based on their observed historical data $\mathbf{r}_{u}$. To measure the difference between the prediction $\hat{\mathbf{r}}_{u}$ and the truth $\mathbf{r}_{u}$, we have the following:
(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}).$
Here, the function $\mathcal{L}_{\text{recon}}$ measures the reconstruction error between $\hat{\mathbf{r}}_{u}$ and $\mathbf{r}_{u}$, and $\alpha_{u}$ represents the loss weight for the client $u$ ($u\in\mathcal{U}$), which is used to balance the client’s contribution to the overall training, and satisfies the condition $\sum_{u\in\mathcal{U}}\alpha_{u}=1$. By optimizing Eq. (1), FedRec aims to make the predicted item scores $\hat{\mathbf{r}}_{u}$ for each user $u$ as close as possible to their true interactions ${\mathbf{r}}_{u}$.
Figure 1. The framework of FedMR. FedMR utilizes a FM to process multimodal data, including visual and textual information, to enhance personalized recommendations. On the server side, FedMR encodes visual and textual data into fixed embeddings. On each client, a Mixing Feature Fusion Module dynamically combines these embeddings with ID embeddings, adjusting the fusion weights based on user interactions to generate personalized item embeddings, capturing fine-grained user diversity.
3.2. ID-based FedRec
Currently, most FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023c, d; Li et al., 2023a, 2024b) construct a global item embedding $\mathbf{D}$ to capture the items’ features and their relationships with neighboring items, forming $\hat{\mathbf{r}}_{u}$ as follows:
(2) $\hat{\mathbf{r}}_{u}=f(\mathbf{D};\theta_{u}),$
where the function $f$, driven by the user-specific parameters $\theta_{u}$, generates personalized predictions $\hat{\mathbf{r}}_{u}$ upon receiving $\mathbf{D}$. However, these methods overlook the rich item multimodal information. Therefore, we consider such methods as ID-based FedRec in this work.
4. Methodology
The aim of our proposed method FedMR is to integrate the multimodal features of items with ID embeddings within a federated environment, and incorporates these features into existing ID-based FedRec methods. By doing so, we transform current FedRec methods into FedMMR methods, allowing them to fully utilize the multimodal features of items to predict items that users have not yet viewed but may find interesting, while protecting user privacy.
4.1. Encoding Multimodal Information with FM
In RecSys, users typically interact with only a small subset of all available items, which are usually stored on the server. Therefore, in multimodal scenarios, we assume the server holds the complete modal information of the item set. FedMR begins by introducing a foundation model $h$ on the server side to encode the items’ modal information, producing feature embeddings for each modality:
(3) $\mathbf{V}=h(\mathbf{E}_{v}),\quad\mathbf{C}=h(\mathbf{E}_{c}).$
Eq. 3 has two main advantages: 1) It transforms the visual or textual modality $\mathbf{E}_{m}$ into a format that the recommendation model can process; 2) Compared to transmitting raw modality, the encoded feature embeddings require less communication overhead, thus improving communication efficiency. Moreover, FedMR also builds an id embedding $\mathbf{D}$ to capture the structure information of items.
4.2. Mixing Feature Fusion Module
Considering the diverse interests and behavior patterns of users, FedMR incorporates a Mixing Feature Fusion Module (MFFM) that adopts multiple modality fusion strategies $\{g_{j}\}$ from $\mathcal{G}$ to fuse the multimodal feature embeddings $\mathbf{V}\text{, }\mathbf{C}$, and the id embedding $\mathbf{D}$ :
(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}),$
where $\gamma_{j}$ is parameters of the fusion strategy $g_{j}$, and $\mathbf{F}_{j}$ is the feature embedding obtained after applying the $j$-th fusion strategy. To capture user diversity by enabling each strategy to focus on specific user groups based on their interaction histories $\mathbf{r}_{u}$, FedMR introduces a user-specific router $\phi_{u}$, driven by the parameters $\varphi_{u}$:
(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}).$
The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$. Then, we have the final fused features $\bar{\mathbf{F}}$ as follows:
(6) $\bar{\mathbf{F}}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
According to Eq. (6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item embeddings $\bar{\mathbf{F}}$, enhancing the system’s ability to capture complex user interaction patterns.
4.3. Objective of FedMR
By replacing $\mathbf{D}$ in Eq. (2) with $\bar{\mathbf{F}}$ obtained from Eq. (6) and then substituting into Eq. (1), we have the objective of FedMR as follows:
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}};\theta_{u}),r_{u,i}).$
Eq. (7) defines the objective of FedMR and is compatible with most existing ID-based FedRec methods. By replacing the original reconstruction error calculation function and the values of $\{\alpha_{u}\}$ in these methods with Eq. (7), they can integrate the multimodal information of items into personalized recommendations while protecting user privacy. This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
4.4. Framework
Fig. 1 illustrates the overall framework of FedMR, which leverages a FM to process multimodal information for enhancing personalized recommendations. On the server side, FedMR uses item cover images as visual modality data and titles as textual modality data. These inputs are processed by the FM to generate fixed visual embeddings $\mathbf{V}$ and text embeddings $\mathbf{C}$. Since both the FM and the input data are fixed, these item feature embeddings do not need to be updated during training. Additionally, the server maintains an ID embedding $\mathbf{D}$ to capture the structural information of the items.
Upon receiving the feature data $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from the server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM. Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion. The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy. The fused features $\bar{\mathbf{F}}$ are then fed into the prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$ for items, and the local model parameters are optimized by minimizing reconstruction loss.
During the entire process, only the ID embedding $\mathbf{D}$ and the gradients of fusion strategies in the MFFM are sent back to the server for aggregation and updates, hereby protecting user privacy while integrating shared information across clients. The core of FedMR lies in dynamically adjusting weights in the MFFM based on the user’s local interactions $\mathbf{r}_{u}$, thereby enhancing the diversity and personalization of the RecSys while ensuring privacy protection. As a result, FedMR is compatible with existing ID-based FedRec methods, allowing multimodal item information to be integrated into these methods. Furthermore, FedMR can be extended by incorporating more modalities or adjusting fusion strategies as needed.
Algorithm 1 FedMR
Input: $\mathbf{R}$, $\eta$, $T$, $A$
Initialize: $\{\theta_{u}\}$, $\{\varphi_{u}\}$, $\{\gamma_{j}\}$, $\mathbf{D}$
GlobalProcedure:
1:  $\mathbf{V}\leftarrow h(\mathbf{E}_{v})$, $\mathbf{C}\leftarrow h(\mathbf{E}_{c})$ according to Eq. (3);
2:  for $t=1,2,\dots,T$ do
3:     $S_{t}$leftarrow randomly select $n_{s}$ from $n$ clients;
4:     for all client index $u\in S_{t}$ do
5:        $\hat{\mathbf{r}}_{u}$, $\mathbf{D}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}$leftarrow ClientUpdate($\varphi$, $\theta$);
6:     end for
7:     $\mathbf{D}\leftarrow\sum_{u\in S_{t}}\alpha_{u}\mathbf{D}^{(u)}$;
8:     $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
9:  end for
10:  return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
ClientUpdate:
1:  $\nabla^{(u)}_{j}{\gamma}\leftarrow 0$, $j=1,2,\cdots|\mathcal{G}|$;
2:  for $a=1,2,\dots,A$ do
3:     Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}$ by differentiating Eq. (7);
4:     Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$;
5:     Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$;
6:     Update $\mathbf{D}\leftarrow\mathbf{D}-\eta\nabla_{\mathbf{D}}$;
7:     Update $\gamma_{j}\leftarrow\gamma_{j}-\eta\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
8:     Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
9:     Compute $\hat{\mathbf{r}}_{u}$ according to Eq. (2);
10:  end for
11:  return: $\hat{\mathbf{r}}_{u}$, $\mathbf{D}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
4.5. Algorithm
To optimize the FedMR objective function defined in Eq. (7), we propose an alternating optimization algorithm for training FedMR. The detailed steps of the algorithm are presented in Alg. 1.
The FedMR framework consists of two main components: the GlobalProcedure and the ClientUpdate. In the GlobalProcedure, the server first uses a FM $h(\cdot)$ to encode the multimodal data, generating the corresponding embeddings $\mathbf{V}$ and $\mathbf{C}$. The server then randomly selects a subset of clients to participate in each training round to reduce the risk of privacy leakage caused by clients participating in consecutive rounds (Li et al., 2023a). For each selected client $u\in S_{t}$, the client update process is executed to obtain the local prediction result $\hat{\mathbf{r}}_{u}$, the local updated id embeddings $\mathbf{D}^{(u)}$, and the local accumulated gradient $\{\nabla^{(u)}_{j}\gamma\}$. The server aggregates the results returned by the clients in a weighted manner, and updates the global shared embedding $\mathbf{D}$ and fusion strategy parameters $\{\gamma_{j}\}$.
During the ClientUpdate, each client calculates the gradients of the model parameters, such as $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$, and $\nabla^{(u)}_{\gamma_{j}}$, based on the FedMR objective function in Eq. (7) and updates these parameters using gradient descent (Amari, 1993). Additionally, each client accumulates the local gradient $\nabla^{(u)}_{\gamma_{j}}$ into $\nabla^{(u)}_{j}\gamma$ for the global update of the fusion strategy parameters. By incorporating a FM $h(\cdot)$ to handle multimodal data $\mathcal{M}$, FedMR achieves efficient personalized multimodal recommendation while preserving user privacy.
4.6. Complexity Analysis
Since FedMR is compatible with various FedRec methods, we focus on the time and space complexity introduced by incorporating multimodal data of items in this work. Assuming $n$ users, $m$ items and $q$ experts, FedMR , where the multimodal information of the items is encoded on the server side into feature embeddings $\mathbf{V}\in\mathbb{R}^{m\times d}$ and $\mathbf{C}\in\mathbb{R}^{m\times d}$ using a foundation model, with $d\ll\{n,m\}$. Since these embeddings remain fixed during training, they contribute a space complexity of $O(md)$ while having negligible impact on time complexity. FedMR also maintains an ID embedding $\mathbf{D}\in\mathbb{R}^{m\times d}$ to capture structural information of the items, with a space complexity of $O(md)$. During each communication round, the server aggregates and updates local data from the clients, requiring a time complexity of $O(\alpha_{u}md(q+1))$, while each client performs the Mixing Feature Fusion Module to fuse features, resulting in a total client-side space complexity of $O(md+K)$ and a time complexity of $O(P)$, where $K$ and $P$ represents the size and cost time of the component parameters, such as Router, MLP, Gate, etc, respectively. Therefore, FedMR requires a total space complexity of $O(n(md+K)+3md)$ and a total time complexity of $O(nmd(q+1)+nP)$.
Table 1. The statistical information of the datasets used in our work includes the following: #Interactions represents the number of observed interactions. #Users denotes the number of users. #Items indicates the number of items.
Dataset #Users #Items #Interactions Sparsity
KU 204 560 3488 96.95%
Dance 10231 1676 80086 99.53%
Food 5990 1125 36482 99.46%
Movie 15908 2528 111091 99.72%
4.7. Privacy Preservation
FedMR uses a distributed architecture based on FL to effectively reduce the risk of user privacy leakage. Except for the Mixing Feature Fusion Module, all other components of FedMR have the same privacy risk level as existing ID-based federated recommendation methods. In the Mixing Feature Fusion Module, only the locally updated ID embeddings $\mathbf{D}^{(u)}$ and the local accumulated gradients ${\nabla^{(u)}_{j}\gamma}$ are sent to the server, without revealing any raw user data, also maintaining the same communication privacy protection strategy as current methods (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024b; Feng et al., 2024; Wu et al., 2024). To protect user privacy, randomly selecting participants in each communication round and avoiding the same participants in consecutive rounds can effectively reduce the risk of privacy leakage (Chai et al., 2020; Li et al., 2023a). Additionally, FedMR can incorporate techniques such as differential privacy (Abadi et al., 2016) and stochastic gradient clipping to further enhance privacy protection. We also conduct an ablation study to evaluate the effectiveness of FedMR when additional privacy-preserving measures are applied.
5. Experiments
5.1. Datasets
In this study, we conducte comprehensive experiments on four downstream datasets from Ninerec [^1] The dataset publicly available on: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a) to evaluate the performance and effectiveness of the proposed method FedMR: KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie). Table 1 provides the detailed statistics of these datasets. The KU dataset is derived from KuaiShou, a well-known short video platform in China, where each entry represents a short video. The Food, Dance, and Movie datasets are collected from the BiliBili platform, focusing on food and cooking videos, dance-related videos, and movie-related videos, respectively. Each dataset presents unique content types and user interaction patterns, but all share a common challenge of high sparsity, with sparsity levels exceeding 95%. For our experiments, we include only users who had interacted with at least 5 items in each dataset to avoid the cold-start problem. Additionally, for missing images in the datasets, we fill them with the average values of visual features; for missing titles, we set ”The title is missing.”
Table 2. Performance comparison of Centralized (Cent) and Federated (Fed) Methods on four real-world datasets using HR@50 and NDCG@50. The federated methods are evaluated in both their original versions (vanilla) and enhanced with FedMR (w/ FedMR). The percentages in parentheses indicate the improvement of the version with FedMR compared to the original version.
KU Dance Food Movie
Method HR@50 NDCG@50 HR@50 NDCG@50 HR@50 NDCG@50 HR@50 NDCG@50
VBPR 0.4412 0.1547 0.2306 0.0747 0.2472 0.0790 0.1429 0.0486
BM3 0.4167 0.1595 0.2654 0.0856 0.2563 0.0820 0.1917 0.0666
Cent MGCN 0.3382 0.1508 0.2363 0.0767 0.2409 0.0777 0.1697 0.0574
vanilla 0.1127 0.0562 0.0782 0.0201 0.0696 0.0191 0.0552 0.0158
FedAvg w/ FedMR 0.5147 (356.70% $\uparrow$) 0.1491 (165.30% $\uparrow$) 0.1007 (28.77% $\uparrow$) 0.0261 (29.85% $\uparrow$) 0.0948 (36.21% $\uparrow$) 0.0268 (40.31% $\uparrow$) 0.0564 (2.17% $\uparrow$) 0.0160 (1.27% $\uparrow$)
vanilla 0.2255 0.0643 0.0676 0.0177 0.0820 0.0219 0.0581 0.0163
FCF w/ FedMR 0.5539 (145.63% $\uparrow$) 0.1597 (148.37% $\uparrow$) 0.0699 (3.40% $\uparrow$) 0.0182 (2.75% $\uparrow$) 0.0918 (11.95% $\uparrow$) 0.0247 (12.79% $\uparrow$) 0.0667 (14.80% $\uparrow$) 0.0180 (10.44% $\uparrow$)
vanilla 0.1176 0.0249 0.0221 0.0055 0.0548 0.0163 0.0185 0.0054
FedNCF w/ FedMR 0.1961 (66.75% $\uparrow$) 0.0609 (144.58% $\uparrow$) 0.0708 (220.36% $\uparrow$) 0.0187 (240.00% $\uparrow$) 0.0843 (53.83% $\uparrow$) 0.0243 (49.08% $\uparrow$) 0.0406 (119.46% $\uparrow$) 0.0125 (131.48% $\uparrow$)
vanilla 0.0637 0.0230 0.0252 0.0068 0.0327 0.0093 0.0170 0.0042
PFedRec w/ FedMR 0.2157 (238.62% $\uparrow$) 0.0616 (167.83% $\uparrow$) 0.0392 (55.56% $\uparrow$) 0.0092 (35.29% $\uparrow$) 0.0836 (155.66% $\uparrow$) 0.0207 (122.58% $\uparrow$) 0.0413 (142.94% $\uparrow$) 0.0132 (214.29% $\uparrow$)
vanilla 0.4706 0.1086 0.0682 0.0186 0.0696 0.0191 0.0453 0.0136
Fed FedRAP w/ FedMR 0.4824 (2.51% $\uparrow$) 0.1305 (20.17% $\uparrow$) 0.0895 (31.19% $\uparrow$) 0.0247 (32.91% $\uparrow$) 0.1142 (64.08% $\uparrow$) 0.0294 (53.93% $\uparrow$) 0.0586 (29.36% $\uparrow$) 0.0199 (46.32% $\uparrow$)
5.2. Baselines
To comprehensively evaluate the performance of FedMR, we select 3 cutting-edge centralized MMRec methods, and 5 ID-based federated methods commonly used in FedRec tasks with implicit feedback to serve as baselines in our study:
VBPR (He and McAuley, 2016) incorporates visual features into personalized ranking to improve recommendation accuracy based on implicit feedback.
BM3 (Zhou et al., 2023a) effectively integrates multimodal features through a self-supervised learning framework.
MGCN (Yu et al., 2023) utilizes a multi-view graph convolutional network to separately capture the multimodal features of users and items.
FedAvg (McMahan et al., 2017) is the earliest federated learning method, performing weighted aggregation based on the relative sizes of local client data.
FCF (Ammad-Ud-Din et al., 2019) introduces collaborative filtering into federated learning, assigning equal weights to all clients for average aggregation.
FedNCF (Perifanis and Efraimidis, 2022) updates user embeddings locally on each client and synchronizes item embeddings on the server for global updates.
PFedRec (Zhang et al., 2023c) aggregates item embeddings from clients on the server side and uses them during evaluation to achieve bilateral personalized recommendations.
FedRAP (Li et al., 2023a) achieves fine-grained personalized modeling of users and items through additive personalization.
The methods VBPR, BM3, and MGCN are three centralized benchmark MMRec methods that demonstrate baseline performances of the used datasets in a centralized recommendation setting. The methods FedAvg, FCF, FedNCF, PFedRec, and FedRAP are five ID-based FedRec methods. We conducted experiments on both the vanilla versions of these ID-based FedRec methods and their multimodal versions integrated with our proposed method FedMR to demonstrate the effectiveness of FedMR. Implementations of all used baseline methods are publicly available through their papers. All methods are implemented using PyTorch (Paszke et al., 2019), and all experiments are run on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, 3 Quadro RTX 6000 GPUs, and 395 GB of RAM.
5.3. Experimental Setting
Following previous studies (Zhang et al., 2023c; Li et al., 2023a), we perform negative sampling for each positive sample in the training set and use the leave-one-out strategy for validation. We conduct hyperparameter tuning for all baseline methods on each dataset used, selecting the learning rate $\eta$ and adjusting the hyperparameters in FedRAP’s objective function from $\{10^{i}\mid i=-4,\ldots,-1\}$. To ensure fairness, we use the CLIP (Radford et al., 2021a) implemented by the Open CLIP (Ilharco et al., 2021; Radford et al., 2021b; Schuhmann et al., 2022) as the foundation model in our experiments, loading the ViT-B-32 parameters trained by OpenAI to process multimodal information. It should be noted that the foundation model in FedMR can be replaced with other models as needed. Additionally, we add a feature mapping layer before the input fusion layer of the multimodal embeddings and ID embeddings, mapping the feature dimensions output by the foundation model to a fixed latent embedding dimension, i.e. 64, to improve storage efficiency. We fix the training batch size at 2048 for all methods, and for methods with hidden layers, the number of layers is set to 3. For each federated baseline, we adopt the same aggregation strategy to update the global part of FedMR. In the main experiments, all clients participated in the training of both the original versions and the FedMR versions of all federated baselines, and no additional privacy protection measures are applied. Additionally, for these federated baselines, the number of local training epochs was set to 10. Except for loading the foundation model, we do not use any pre-training strategies for any methods.
5.4. Evaluation
In our experiments, we treat the observed interactions as positive samples and the rest as negative samples. We sort the predicted interaction scores between users and all candidate items in descending order, masking the observed positive samples from the training set. Then, we evaluate the recommendation performance using two widely-used metrics: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020). In the experiments, we set $K=50$.
5.5. Comparison Analysis
Table 2 shows the performance of both centralized and federated methods across the four datasets. The results highlight that FedMR significantly boosts the performance of the used FedRec methods, particularly in the KU dataset, where HR@50 for FedAvg, FCF, PFedRec, and FedRAP exhibits notable improvements. This indicates that FedMR effectively enhances traditional ID-based FedRec methods by integrating multimodal item information. The performance improvements in other datasets further demonstrate FedMR’s ability to alleviate data sparsity by better leveraging users’ interaction histories alongside multimodal data for more personalized recommendations, while protecting user privacy. Compared to centralized MMRec methods, FedRec approaches using FedMR achieve similar or even superior performance while maintaining privacy, e.g., in the KU dataset, FedAvg with FedMR surpasses all centralized methods in terms of HR@50. Federated methods with FedMR also perform strongly across other datasets, reinforcing FedMR’s capability to improve recommendation quality without sacrificing user privacy.
Table 2 further shows that by incorporating multimodal information, FedMR provides a more comprehensive representation of item features. Compared to ID-based FedRec, the use of multimodal data enables more accurate predictions of user preferences. Additionally, the MFFM of FedMR dynamically adjusts the weights of visual, textual, and ID features based on each user’s dependence on these modalities. This flexible approach to feature fusion avoids the drawbacks of single fusion strategy methods, allowing the model to better capture complex relationships between modalities and further enhance personalized recommendation accuracy.
5.6. Ablation Study
To validate the effectiveness of each component in FedMR, we conducted experiments using several variants of the system. Specifically, we independently applied the Sum, MLP, and Gate strategies for the fusion of multimodal features and ID features, to assess the impact of dynamic weighted fusion based on users’ interaction history $\mathbf{r}_{u}$, $\forall u\in\mathcal{U}$. Additionally, we introduce random Gaussian noise into the gradients uploaded to the server to evaluate the potential performance impact of this privacy-preserving measure. We also examine the distribution of multimodal and ID features within the same feature space. These ablation studies demonstrate the effectiveness of each component of FedMR.
Figure 2. Visualization of multimodal features and ID features mapped into a shared representation space.
5.6.1. Ablation Study on Modality Distributions
To investigate the distribution relationship between multimodal features and ID features, we select FedRAP (Li et al., 2023a) as the backbone for FedMR in this study. FedRAP maintains both globally shared and user-specific ID features, which helps demonstrate how different ID features affect the distribution within the feature space.
Fig. 2 visualizes the distribution of modality features and ID features after training FedRAP with FedMR, on the KU dataset. The first subfigure in Fig. 2 shows that when the multimodal features of items and globally shared ID features are mapped into a shared representation space, they remain in distinct subregions. This phenomenon, known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a), helps the model distinguish between different modalities, reducing noise interference and improving the robustness of multimodal data processing. The following two subfigures incorporate the personalized ID embeddings of the users 00 and $42$, respectively, highlighting the significant impact of personalized ID features on the distribution in the shared space. By comparing the first subfigure with the subsequent two, we observe that the introduction of personalized ID features causes the ID features to gradually expand towards the regions of visual and textual features, shrinking the modality gap. This facilitates the fusion of multimodal information and enhances the accuracy of personalized recommendations. Furthermore, by comparing the differences in modality gap changes for each user, it becomes evident that the model can adjust the feature distribution based on user-specific behaviors and preferences, thus better satisfying individual user needs.
This study underscores the importance of personalized ID features in federated recommendation systems. Due to privacy constraints, the model cannot access raw user data directly, but through the use of personalized ID features, the model can provide personalized recommendation services while maintaining user privacy.
Table 3. Performance comparison of FedRec Methods on HR@50 and NDCG@50 with and without multimodal and ID features on the KU dataset.
Method Metric FedMR w/o $\mathbf{V}$ w/o $\mathbf{C}$ w/o $\mathbf{D}$
FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.5094 (1.03% $\downarrow$)
NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1407 (5.66% $\downarrow$)
FCF HR@50 0.5539 0.1569 (71.67% $\downarrow$) 0.1569 (71.67% $\downarrow$) 0.1275 (76.99% $\downarrow$)
NDCG@50 0.1597 0.0486 (69.57% $\downarrow$) 0.0476 (70.19% $\downarrow$) 0.0287 (82.04% $\downarrow$)
FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1602 (18.31% $\downarrow$)
NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0538 (11.66% $\downarrow$)
PFedRec HR@50 0.2157 0.2010 (6.82% $\downarrow$) 0.1225 (43.21% $\downarrow$) 0.0490 (77.27% $\downarrow$)
NDCG@50 0.0616 0.0557 (9.58% $\downarrow$) 0.0341 (44.64% $\downarrow$) 0.0191 (68.97% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2647 (30.78% $\downarrow$) 0.1863 (51.29% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0734 (43.75% $\downarrow$) 0.0490 (62.43% $\downarrow$)
5.6.2. Ablation Study on Modality Efficiency
Table 3 presents a performance comparison of the five used ID-based FedRec methods with the inclusion of multimodal information. It evaluates these methods under the full model (FedMR) and with the removal of visual modality (w/o V), textual modality (w/o C), and ID features (w/o D), focusing on HR@50 and NDCG@50 metrics.
Firstly, removing the visual modality results in a noticeable performance drop across all methods, particularly for FCF (HR@50 decreased by 71.67%) and FedRAP (HR@50 decreased by 30.78%). This highlights the critical role of visual information in RecSys, especially for FCF. Visual features provide rich, intuitive image-based information that helps models better capture user preferences regarding the appearance of items. Without this visual context, models struggle to accurately infer user preferences, leading to a significant decline in recommendation performance. Secondly, removing the ID feature also causes a substantial decline in performance. The ID feature is essential for personalized recommendations, as it captures users’ historical interactions and reflects their past behaviors and interests. When the ID feature is removed, the model loses its primary mechanism for personalization, resulting in decreased accuracy and personalization of recommendations. This impact is especially pronounced in methods like PFedRec (HR@50 decreased by 82.82%) and FedRAP (HR@50 decreased by 51.92%), which heavily rely on ID features. In contrast, removing the text modality has a relatively smaller impact, though it still affects some methods. For instance, in FedAvg and PFedRec, removing the text modality led to HR@50 decreases of 27.63% and 43.21%, respectively. This indicates that text information plays an important role in capturing semantic relationships between user behaviors and item descriptions, helping the model better understand user needs and item characteristics, thereby enhancing recommendation performance.
In summary, all methods experienced performance degradation to varying degrees when any modality was removed. Considering Fig. 2, visual, text, and ID features are often complementary. Removing any one of these disrupts the balance, making it difficult for the model to fully represent user preferences and item characteristics. Additionally, different methods depend on different modalities to varying extents. For instance, FCF and PFedRec are more dependent on visual information, whereas FedAvg is less reliant on ID features. These differences in dependency lead to varying levels of performance decline when specific modalities are removed. Further experimental results are provided in the appendix.
Figure 3. Performance comparison of different feature fusion strategies in various FedRec methods combined with FedMR on the KU dataset.
5.6.3. Ablation Study on Fusion Strategy
Fig. 3 presents the HR@50 and NDCG@50 evaluation results for five used FedRec methods using different feature fusion strategies with FedMR on the KU dataset. The results indicate that different fusion strategies significantly impact performance. Notably, the Mix strategy, which dynamically combines multiple fusion methods based on user interaction history, outperforms others, while the Sum strategy shows poor performance. This trend is consistent across all metrics.
The Sum strategy simply adds different modality features without considering their complex interactions and inter-relationships. Since each modality’s importance varies, simple addition can lead to information loss or imbalanced weighting, making it difficult to capture relationships effectively. Additionally, RecSys often face a long-tail distribution, where a few popular items dominate user interactions (Park and Tuzhilin, 2008; Yin et al., 2012). The simplicity of the Sum strategy may lead to overfitting on popular items, resulting in poor performance on less interacted items and affecting overall model performance.
In contrast, the MLP strategy uses fully connected networks to learn nonlinear modality combinations, which better captures complex interactions between item modalities and ID features. However, its effectiveness is limited by network depth and the number of parameters. On the other hand, the Gate strategy learns adaptive weights to dynamically filter the contribution of each modality, optimizing the importance of different features.
The Mix strategy integrates all fusion methods and adjusts their weights dynamically based on user interaction history, allowing the model to capture more fine-grained personalized item representations. This approach enables more accurate matching of user preferences and effectively captures complex relationships between multimodal features, leading to superior performance.
This study highlights the importance of personalized item embeddings in FedRec. In FedMR, each client dynamically adjusts fusion strategy weights through the Mixing Feature Fusion Module, based on user interaction history, to generate personalized item embeddings, enhancing personalization and accuracy while maintaining user privacy, thus improving overall system performance.
Table 4. Performance comparison of FedRec methods combined with FedMR, with and without Gaussian noise (mean=0, variance=0.1) on KU dataset for privacy protection.
Method Metric FedMR w/ noise Degrade
FedAvg HR@50 0.5147 0.4969 3.46% $\downarrow$
NDCG@50 0.1491 0.1416 5.03% $\downarrow$
FCF HR@50 0.5539 0.5378 2.91% $\downarrow$
NDCG@50 0.1597 0.1402 12.21% $\downarrow$
FedNCF HR@50 0.1961 0.1680 14.33% $\downarrow$
NDCG@50 0.0609 0.0420 31.03% $\downarrow$
PFedRec HR@50 0.2157 0.1965 8.90% $\downarrow$
NDCG@50 0.0616 0.0586 4.87% $\downarrow$
FedRAP HR@50 0.3824 0.3725 2.59% $\downarrow$
NDCG@50 0.1305 0.1236 5.29% $\downarrow$
5.6.4. Privacy Protection
Table 4 presents the experimental results of the used FedRec methods combined with FedMR, comparing the performance with and without adding Gaussian noise (mean=0, variance=0.1) to the gradients of the global components uploaded to the server, to protect user privacy. The experiments are conducted on the KU dataset during feature fusion. The results indicate that all methods experienced some performance degradation after adding noise, particularly on the NDCG metric, which is more sensitive to ranking accuracy. This is because noise disturbs the gradient precision, which affects the model’s ability to update effectively.
Among the methods, FedNCF is the most sensitive to noise, with HR@50 dropping by 14.33% and NDCG@50 by 31.03%. As a complex neural network-based model, FedNCF depends on numerous parameters to capture nonlinear features. The introduction of noise disrupts these feature interactions, resulting in a significant decline in performance. In contrast, FedAvg and FedRAP show greater robustness to noise, with only minimal performance degradation. This suggests that these methods are more effective at maintaining recommendation accuracy while ensuring privacy protection.
In conclusion, although adding noise helps protect user privacy, its impact varies based on model complexity and feature interaction mechanisms. Simpler models like FedAvg and adaptive models like FedRAP are more resilient to noise, while complex neural network models like FedNCF experience more significant performance drops due to the noise interference.
6. Conclusion
This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client. This module dynamically adjusts fusion weights based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences, thereby enhancing personalized recommendations. FedMR allows flexible modification of fusion strategies and supports different modalities of item data as needed. Experimental results show that FedMR performs well on four multimodal recommendation datasets and is compatible with existing ID-based methods without requiring changes to the underlying framework. This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.
Appendix A Algorithm Optimization
A.1. Derivation of the Update for $\gamma_{j}^{(u)}$
The update process for $\{\gamma_{j}^{(u)}\}_{j=1}^{|\mathcal{G}|}$ in Alg. 1 is based on gradient descent. The local update at each client $u$ during the iteration $t$ follows the following rule:
(8) $\gamma_{j}^{(t)}=\gamma_{j}^{(t-1)}-\eta\nabla_{j}^{(u)}\gamma^{(t)},$
where $\eta$ is the learning rate, and $\nabla_{j}^{(u)}\gamma^{(t)}$ represents the gradient of $\gamma_{j}^{(u)}$ at the iteration $t$ on client $u$.
To capture how $\gamma_{j}^{(u)}$ evolves over time, we can unroll this equation over several iterations. By recursively applying the update rule in Eq. (8), we obtain the following:
(9) $\gamma_{j}^{(t)}=\gamma_{j}^{(0)}-\eta\sum_{i=1}^{t}\nabla_{j}^{(u)}\gamma^{(i)},$
where $\gamma_{j}^{(0)}$ is the initial value of $\gamma_{j}^{(u)}$, and the second term is the sum of all gradients computed up to the $t$-th iteration. This expression shows that the updated value of $\gamma_{j}^{(u)}$ depends on its initial value and the accumulated gradient over $t$ iterations.
A.2. Accumulated Gradient
The accumulated gradient $\nabla_{j}^{(u)}\gamma$ at the client $u$ is the sum of the individual gradients across the iterations, which be expressed as:
(10) $\nabla_{j}^{(u)}\gamma=\frac{1}{\eta}\left(\gamma_{j}^{(0)}-\gamma_{j}^{(t)}\right).$
Eq. (10) represents the total change in $\gamma_{j}^{(u)}$ during the local training process. It relates the initial value $\gamma_{j}^{(0)}$ to the current value $\gamma_{j}^{(t)}$, scaled by the learning rate $\eta$.
A.3. Global Aggregation
In federated learning, the global model is updated by aggregating the local updates from multiple clients. For the parameter $\gamma_{j}$, this is achieved by summing the accumulated gradients $\nabla_{j}^{(u)}\gamma$ from all selected clients. The global update for $\gamma_{j}$ can be written as:
(11) $\gamma_{j}=\gamma_{j}-\eta\sum_{u\in\mathcal{S}_{t}}\alpha_{u}\nabla_{j}^{(u)}\gamma,$
where $\mathcal{S}_{t}$ is the set of clients selected during round $t$. This step updates the global parameter $\gamma_{j}$ by aggregating the contributions from each client’s local updates. The derived equations demonstrate that the update process is consistent with the principles of gradient-based optimization. This ensures that $\gamma_{j}^{(u)}$ is updated in an efficient and mathematically sound manner within the FL framework.
Table 5. Changes of parameter numbers in FedRec with and without FedMR on the KU Dataset, Compared to Centralized Recommendation Systems.
Method KU Dance Food Movie
Cent VBPR 127552 1482432 904320 2263616
BM3 1752576 8600576 5582848 12815872
MGCN 708864 2564800 1693888 3855104
Fed FedAvg Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FCF Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FedNCF Vanilla 108737 1535041 921665 1182689
FedMR 564292 2561988 1666500 3833924
PFedRec Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FedRAP Vanilla 71745 214593 144065 323649
FedMR 742340 1885124 1320900 2757572
Appendix B Parameter Analysis
Table 5 illustrates the change in the number of learnable parameters on a single client in the KU dataset when FedMR is integrated into FedRec or not, compared to centralized recommendation systems. As shown in the table, the introduction of FedMR significantly increases the parameter size. This is because FedMR incorporates multimodal information (e.g., visual and textual features) alongside ID embeddings, increasing the model’s complexity and consequently requiring more parameters to learn.
Despite the increase in the number of learnable parameters, Table 2 demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset. By combining multimodal features with ID embeddings, FedMR better leverages users’ interaction histories to capture the fine-grained personality of users. Additionally, the mixing feature fusion module in FedMR adjusts the weights of multimodal features based on users’ interaction history, greatly enhancing both the accuracy and personalization of recommendations. Moreover, compared to traditional FedRec methods, which rely solely on the ID embeddings, FedMR enables personalized modeling by integrating multimodal information, thereby further improving recommendation quality. While the model’s complexity increases, this trade-off results in significant performance gains.
As privacy protection is crucial in the federated settings, FedMR processes multimodal data locally on the client and only uploads essential parameters, e.g., ID embeddings and gradients, to the server, avoiding the exposure of raw user data. Although this adds computational overhead and increases parameter size on the client, these sacrifices are necessary to ensure privacy while also boosting recommendation performance. Thus, the increased learnable parameter size is not only a necessary cost for improved recommendation performance but also a vital measure for ensuring privacy.
Appendix C More Experiments
Figure 4. Performance comparsions of different FedRec methods with FedMR under varying client sampling ratios on HR@50 and NDCG@50.
C.1. Ablation Study on Client Sampling
To investigate the impact of introducing FedMR for multimodal data fusion on different FedRec methods under varying client sampling ratios, we conduct experiments on the five FedRec methods using the KU dataset. Fig. 4 presents the results for HR@50 and NDCG@50 as the client sampling ratio changes. Overall, as the client sampling ratio increases, most methods show improvements in HR@50 and NDCG@50, indicating that more client participation helps enhance the personalization of recommendations. However, the performance improvement varies across different methods. For PFedRec, the increase in HR@50 and NDCG@50 after introducing FedMR is relatively small, possibly due to its lower sensitivity to the number of participating clients. This suggests that PFedRec may rely more on the quality of individual client data rather than the number of clients. As a result, the advantages of FedMR’s multimodal fusion are not fully realized when the client count is low. Additionally, when the client sampling ratio is low, the model’s performance becomes unstable, indicating that insufficient data affects the stability of feature learning. This is especially true in multimodal scenarios, where a lack of sufficient modality information weakens the fusion effect, reducing recommendation accuracy.
Appendix D Limitations and Future Work
While FedMR has shown promising improvements in federated multimodal recommendation systems, several limitations remain:
  1. (1)
    Data Heterogeneity: Considering the federated settings, the distribution of user data varies significantly. Although FedMR mitigates this issue by integrating multimodal information, its performance is still limited when handling highly heterogeneous data. Future work could focus on introducing more robust personalization models or adaptive aggregation strategies to better address the differences between users.
  2. (2)
    Communication Overhead: Although FedMR reduces some communication costs by encoding multimodal data, processing multimodal features still results in high computational and communication costs. Future research could explore optimizing communication protocols or using model compression techniques to further minimize these costs.
  3. (3)
    Limited Flexibility in Multimodal Fusion: While FedMR supports dynamic adjustments to its fusion strategies, its current approach may lack flexibility when dealing with complex user behaviors. Future work could explore more adaptive and layered fusion methods to better capture the diverse preferences of users.
In conclusion, future developments could enhance the overall performance of federated recommendation systems by improving the handling of data heterogeneity, reducing communication costs, and optimizing multimodal fusion strategies.

### v2

@@
-Personalized Item Embeddings in Federated Multimodal Recommendation
+Personalized Item Representations in Federated Multimodal Recommendation
 Abstract.
-Federated recommendation systems play a crucial role in protecting user privacy.
-However, existing methods primarily rely on ID-based item embeddings, overlooking the rich multimodal information of items.
-To address this limitation, we propose a novel Federated Multimodal Recommendation System called FedMR.
-FedMR leverages a foundation model on the server side to encode multimodal data, such as images and text, associated with items.
-To tackle the challenge of data heterogeneity caused by varying user preferences, FedMR introduces a Mixing Feature Fusion Module on the client.
-This module dynamically adjusts the weights of different fusion strategies based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences.
-FedMR is compatible with existing ID-based federated recommendation systems, improving their performances without modifying the original framework.
-Our experiments on four real-world multimodal recommendation datasets demonstrate the effectiveness of FedMR.
-Our code is available at https://anonymous.4open.science/r/FedMR.
+Federated recommendation systems are essential for providing personalized recommendations while protecting user privacy.
+However, current methods mainly rely on ID-based item embeddings, neglecting the rich multimodal information of items.
+To address this, we propose a Federated Multimodal Recommendation System, called FedMR.
+FedMR uses a foundation model on the server to encode multimodal item data, such as images and text.
+To handle data heterogeneity caused by user preference differences, FedMR introduces a Mixing Feature Fusion Module on each client, which adjusts fusion strategy weights based on user interaction history to generate personalized item representations that capture users’ fine-grained preferences.
+FedMR is compatible with existing ID-based federated recommendation systems, improving performance without modifying the original framework.
+Experiments on four real-world multimodal datasets demonstrate FedMR’s effectiveness.
+The code is available at https://anonymous.4open.science/r/FedMR.
 Federated Recommendation, Multimodal Learning, Personalization
@@
 Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity.
-Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item feature representations can significantly improve recommendation results.
+Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item representations can significantly improve recommendation results.
 Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data.
-Thus, generating personalized item embeddings using multimodal data while preserving user privacy remains a significant challenge.
+Thus, generating personalized item representations from multimodal data while ensuring user privacy remains a significant challenge.
 As a potential solution, Multimodal Recommendation Systems (MMRec) integrate data from different modalities, significantly enhancing recommendation performance (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a).
@@
 By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024).
-FedMMRec is still in the early stages of research, and related studies are limited.
+To the best of our knowledge, FedMMRec is still in the early stages of research, and related studies are limited.
 FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering.
@@
 Main Contribution.
-To integrate multimodal item information into FedRec while ensuring user privacy, we propose a Federated Multimodal Recommendation system (FedMR) based on foundation models.
-FedMR uses a foundation model on the server side to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items.
-On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts fusion strategies based on user interaction history, combining multimodal and ID embeddings to generate personalized item embeddings.
-This approach provides more personalized recommendations.
-Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients.
+To integrate multimodal item information into FedRec while ensuring user privacy, we propose a novel Federated Multimodal Recommendation system (FedMR).
+FedMR uses a foundation model on the server to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items.
+On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts the importances of fusion strategies based on user interaction history, combining multimodal and ID embeddings to model fine-grained personalized item representations for users.
+Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients, providing more personalized recommendations.
 The main contributions of this work are summarized as follows:
   • We propose FedMR, a federated multimodal recommendation system using foundation models, which efficiently leverages multimodal data while preserving user privacy;
-  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item embeddings, enhancing personalization and flexibility in recommendations;
+  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item representations, enhancing personalization and flexibility in recommendations;
   • FedMR is compatible with existing ID-based federated recommendation methods, improving recommendation performance without modifying the existing framework;
@@
 In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d).
-However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely heavily on ID embeddings to model items, overlooking the rich multimodal information (e.g., titles and covers) that items may have.
+However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have.
 AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec.
@@
 FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network.
-While these methods achieve personalized item embeddings, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
+While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
 2.2. Multimodal Recommendation Systems
@@
 The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$.
-Then, we have the final fused features $\bar{\mathbf{F}}$ as follows:
-(6) $\bar{\mathbf{F}}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
+Then, we have the final fused features $\bar{\mathbf{F}}^{(u)}$ as follows:
+(6) $\bar{\mathbf{F}}^{(u)}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
 According to Eq.
-(6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item embeddings $\bar{\mathbf{F}}$, enhancing the system’s ability to capture complex user interaction patterns.
+(6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item representations $\bar{\mathbf{F}}^{(u)}$, thereby enhancing the system’s ability to capture fine-grained user preferences.
 4.3. Objective of FedMR
 By replacing $\mathbf{D}$ in Eq. (2) with $\bar{\mathbf{F}}$ obtained from Eq. (6) and then substituting into Eq. (1), we have the objective of FedMR as follows:
-(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}};\theta_{u}),r_{u,i}).$
+(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}^{(u)};\theta_{u}),r_{u,i}).$
 Eq.
@@
 Additionally, the server maintains an ID embedding $\mathbf{D}$ to capture the structural information of the items.
-Upon receiving the feature data $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from the server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM.
+Upon receiving the features $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM.
 Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion.
 The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy.
-The fused features $\bar{\mathbf{F}}$ are then fed into the prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$ for items, and the local model parameters are optimized by minimizing reconstruction loss.
+The personalized item representations $\bar{\mathbf{F}}^{(u)}$ are then fed into a prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$, and optimize the model parameters by minimizing reconstruction loss in Eq.
+(7).
 During the entire process, only the ID embedding $\mathbf{D}$ and the gradients of fusion strategies in the MFFM are sent back to the server for aggregation and updates, hereby protecting user privacy while integrating shared information across clients.
@@
 This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client.
-This module dynamically adjusts fusion weights based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences, thereby enhancing personalized recommendations.
-FedMR allows flexible modification of fusion strategies and supports different modalities of item data as needed.
-Experimental results show that FedMR performs well on four multimodal recommendation datasets and is compatible with existing ID-based methods without requiring changes to the underlying framework.
+This module dynamically adjusts fusion weights based on user interaction history, generating personalized item representations that capture fine-grained user preferences, thereby enhancing personalized recommendations.
+It allows flexible modification of fusion strategies and supports different modalities of item data as needed.
+Experimental results on four multimodal recommendation datasets show that FedMR performs well and is compatible with existing ID-based methods without requiring changes to the underlying framework.
 This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.

### v3

@@
-Personalized Item Representations in Federated Multimodal Recommendation
+Dynamic Fusion Strategies for Federated Multimodal Recommendations
 Abstract.
-Federated recommendation systems are essential for providing personalized recommendations while protecting user privacy.
-However, current methods mainly rely on ID-based item embeddings, neglecting the rich multimodal information of items.
-To address this, we propose a Federated Multimodal Recommendation System, called FedMR.
-FedMR uses a foundation model on the server to encode multimodal item data, such as images and text.
-To handle data heterogeneity caused by user preference differences, FedMR introduces a Mixing Feature Fusion Module on each client, which adjusts fusion strategy weights based on user interaction history to generate personalized item representations that capture users’ fine-grained preferences.
-FedMR is compatible with existing ID-based federated recommendation systems, improving performance without modifying the original framework.
-Experiments on four real-world multimodal datasets demonstrate FedMR’s effectiveness.
-The code is available at https://anonymous.4open.science/r/FedMR.
+Delivering deeply personalized recommendations necessitates understanding user interactions with diverse multimedia features, but achieving this within the constraints of Federated Recommendation Systems (FedRec) is severely hampered by communication bottlenecks, user heterogeneity, and the complexity of privacy-preserving multimodal fusion.
+To this end, we propose FedMR, a novel multimodal FedRec framework centered around the Mixing Feature Fusion Module (MFFM).
+FedMR employs a two-stage process: (1) Server-side centralized multimedia content processing provides rich, shared item context using pre-trained models, mitigating limitations from client sparsity and resource constraints efficiently.
+(2) Client-Side Personalized Refinement, where the MFFM dynamically adapts these server-provided multimodal representations based on client-specific interaction patterns, effectively tailoring recommendations and resolving heterogeneity in user preferences towards different modalities.
+Extensive experiments validate that FedMR seamlessly enhances existing ID-based FedRecs, effectively transforming them into high-performing federated multimodal systems.
 Federated Recommendation, Multimodal Learning, Personalization
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
 1. Introduction
-Under privacy protection regulations, such as GDPR (Voigt and Von dem Bussche, 2017), that emphasize user data privacy, Federated Recommendation Systems (FedRec) have become an essential technology in modern life (Sun et al., 2024).
-However, in these systems, since each client typically only has access to a single user’s interaction history, data sparsity is a common issue that degrades recommendation performance (Li and Long, 2024).
-Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity.
-Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item representations can significantly improve recommendation results.
-Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data.
-Thus, generating personalized item representations from multimodal data while ensuring user privacy remains a significant challenge.
-As a potential solution, Multimodal Recommendation Systems (MMRec) integrate data from different modalities, significantly enhancing recommendation performance (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a).
-Unlike single-modal systems, MMRec can capture hidden relationships and complementary information between various modalities, leading to a better understanding of user preferences and item characteristics (Zhou et al., 2023b).
-For example, MMRec methods (Wei et al., 2019; Wang et al., 2021; Tao et al., 2022) combine video frames, audio tracks, and item descriptions when users interact with online platforms, providing a more accurate representation of user preferences.
-To effectively handle multimodal data, MMRec often uses Foundation Models (FM) pretrained on large-scale datasets, which possess powerful feature extraction and representation capabilities (Bommasani et al., 2021).
-By applying FMs, MMRec can efficiently extract and integrate information from multiple modalities, capturing commonalities and complementary relationships between them (Radford et al., 2021a; Cui et al., 2022; Geng et al., 2022, 2023; Wu et al., 2024).
-However, these centralized MMRec methods rely on servers to aggregate user data, digital activities, and preferences to train accurate models (Liu et al., 2024), which raises privacy concerns (Chai et al., 2020; Li et al., 2023a).
-To address these challenges, Federated Multimodal Recommendation Systems (FedMMRec) have emerged, which integrates MMRec and introduces the capability of processing multimodal information into FedRec.
-By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024).
-To the best of our knowledge, FedMMRec is still in the early stages of research, and related studies are limited.
-FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering.
-AMMFRS (Feng et al., 2024) introduces an attention mechanism on the client side to perform weighted fusion of multimodal data, effectively distinguishing and utilizing the contributions of different modalities.
-These methods not only protect user privacy but also effectively integrate multimodal features to improve recommendation accuracy and personalization.
-However, they rely on fixed multimodal fusion strategies, neglecting the personalized representation of item features on clients, which limits their ability to adapt to diverse user needs.
-Although AMMFRS uses an attention mechanism to distinguish the importance of different modalities, it lacks adaptability to the diversity of user data, resulting in limited flexibility and accuracy of fusion strategies (Niu et al., 2021).
-Main Contribution.
-To integrate multimodal item information into FedRec while ensuring user privacy, we propose a novel Federated Multimodal Recommendation system (FedMR).
-FedMR uses a foundation model on the server to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items.
-On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts the importances of fusion strategies based on user interaction history, combining multimodal and ID embeddings to model fine-grained personalized item representations for users.
-Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients, providing more personalized recommendations.
-The main contributions of this work are summarized as follows:
-  • We propose FedMR, a federated multimodal recommendation system using foundation models, which efficiently leverages multimodal data while preserving user privacy;
-  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item representations, enhancing personalization and flexibility in recommendations;
-  • FedMR is compatible with existing ID-based federated recommendation methods, improving recommendation performance without modifying the existing framework;
-  • Extensive experiments are conducted on four real-world multimodal recommendation datasets to validate the effectiveness of our proposed method FedMR;
-  • We provide the source code to support the reproducibility.
+Recommendation Systems (RecSys) are integral components of modern digital platforms, guiding users through vast item catalogs by suggesting relevant content (Li and Long, 2024).
+Traditional RecSyss rely heavily on collaborative filtering signals derived from user-item interaction histories (Chen et al., 2017).
+However, the rich information embedded within item multimedia content, such as product images, video trailers, textual descriptions, and audio samples, offers a powerful complementary signal (Deldjoo et al., 2020; Zhou et al., 2023b).
+Leveraging these multimodal cues allows Multimodal Recommendation Systems (MMRec) (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a; Yu et al., 2023) to gain a deeper understanding of item characteristics and user preferences, capture subtle nuances missed by interaction data alone, significantly alleviate the cold-start problem for new items (Pan et al., 2022), and ultimately enhance recommendation accuracy and user experience.
+Despite their effectiveness, centralized MMRec approaches typically require aggregating vast amounts of user interaction data and potentially sensitive item features on central servers (Liu et al., 2024).
+This practice raises significant privacy concerns, particularly with the enforcement of stringent data protection regulations like the GDPR (Voigt and Von dem Bussche, 2017).
+To address these concerns, Federated Learning (FL) (McMahan et al., 2017) has emerged as a compelling paradigm, enabling collaborative model training across distributed clients, e.g., user devices, without exposing raw local data.
+Federated Recommendation Systems (FedRec) (Yang et al., 2020; Wang et al., 2022) adapt FL principles to the recommendation domain, offering a promising path towards privacy-preserving personalization.
+However, integrating rich multimedia information into the FL setting poses distinct challenges beyond traditional ID-based FedRecs (Feng et al., 2024).
+First, processing raw, high-dimensional multimedia data on resource-constrained client devices is often computationally prohibitive, hindering feature extraction and user participation (Li et al., 2024a).
+Second, transmitting large multimedia features incurs significant communication overhead and raises privacy concerns, as shared features might reveal sensitive information (Chai et al., 2020; Li et al., 2023a).
+Third, data heterogeneity (Yuan et al., 2024; Li et al., 2019) extends to user preferences for different modalities, such as visual vs.
+textual (Liu et al., 2019; Niu et al., 2021).
+This makes a single, global fusion strategy suboptimal for personalized multimodal understanding (Feng et al., 2024; Li et al., 2024a).
+Finally, client-side data sparsity (Sun et al., 2024) limits item understanding, especially for cold-start items (Zhang et al., 2024c).
+While multimedia data could alleviate this (He and McAuley, 2016; Zhou et al., 2023a), its utilization is obstructed by the aforementioned constraints.
+Hence, designing a FedRec that integrate multimodal data while balancing efficiency, privacy, heterogeneity, and sparsity is thus a critical research problem (Feng et al., 2024; Li et al., 2024a).
+To address these interconnected challenges, we propose FedMR, a novel Federated Multimodal Recommendation framework.
+FedMR utilizes a collaborative server-client architecture where the server leverages powerful pre-trained models for centralized, efficient encoding of item modalities, creating rich, shared contextual representations that combat client-side item understanding limitations due to sparsity.
+Crucially, FedMR introduces the Mixing Feature Fusion Module (MFFM), operating on the client-side for dynamic and personalized adaptation.
+The MFFM empowers each client to learn user-specific weights for combining different fusion perspectives of the server-provided multimodal representations and collaborative signals from ID embeddings, guided by the client’s unique interaction history.
+This adaptive multimodal fusion directly tackles data heterogeneity by tailoring the final item representation to individual user preferences and local data characteristics, leading to more accurate and personalized recommendations while preserving privacy.
+Our main contributions are summarized as follows:
+  • Framework Innovation: We propose FedMR, a novel federated architecture employing centralized server-side processing of item modalities to enrich item context and alleviate client-side limitations while protecting user privacy.
+  • Dynamic Feature Fusion: We introduce the Mixing Feature Fusion Module for dynamic, personalized adaptation of multimodal representations, effectively addressing data heterogeneity and diverse modality preferences.
+  • Seamless Compatibility and Enhancement: FedMR seamlessly integrates with existing ID-based FedRec frameworks, upgrading them with multimodal capabilities while requiring minimal modification and boosting performance.
+  • Comprehensive Empirical Validation: Extensive experiments on real-world multimodal datasets validate FedMR’s effectiveness, showing significant improvements over FedRec baselines.
+We also provide code for reproducibility.
 2. Related Work
 2.1. Federated Recommendation Systems
-Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020).
-In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d).
-However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have.
-AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec.
-AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features.
-In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy.
-Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks.
-Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features.
-FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network.
-While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
+FedRecs aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020).
+Significant progress has been made in FedRec across various tasks, including rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024b), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and cold-start scenarios (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d; Li et al., [n.
+d.]).
+However, many existing FedRec approaches (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) primarily use ID embeddings for item representations, thereby neglecting rich multimodal information like associated text and images.
+Recent efforts have started applying multimodal learning (Baltrušaitis et al., 2018) to FedRec, called Federated Multimodal Recommendation System (FedMMRec).
+AMMFRS (Feng et al., 2024) employs an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data based on feature importance.
+In contrast, FedMMR (Li et al., 2024a) offloads heavy multimodal computation to the server and employs secure cross-client convolution with secret sharing to capture high-order interactions privately.
+Other works focus on enhancing personalization within FedRec.
+PFedRec (Zhang et al., 2023c) introduced dual personalization for both users and items.
+FedRAP (Li et al., 2023a) refined this with additive personalization.
+FedDAE (Li et al., 2024c) uses a variational autoencoder and a gating network to adaptively fuse shared and personalized item knowledge.
+Furthermore, recent works explore integrating Foundation Models (FMs) directly within FedRec.
+FedPA (Zhang et al., 2024c) introduces personalized adapters for lightweight, client-specific FM fine-tuning.
+TransFR (Zhang et al., 2024b) uses FMs to generate universal textual item representations, replacing traditional IDs to improve adaptability and address cold-start issues.
+Despite these advancements in personalization and FM integration, methods predominantly relying on ID embeddings or single-modality representations still limit the full utilization of rich, multi-faceted multimodal item information within the federated setting (Che et al., 2023).
 2.2. Multimodal Recommendation Systems
-Unlike traditional Recommendation Systems (RecSys) that rely primarily on user-item interaction data (such as clicks or ratings) to capture user behavior (Yu et al., 2023), Multimodal Recommendation systems (MMRec) aim to improve the accuracy of recommendations and enhance user experience by utilizing multiple modalities of data , such as text, images, audio, and video (Zhou et al., 2023b).
-As unique identifiers for items, item IDs not only contain semantic information about the items but also capture the characteristics of neighboring items (Yuan et al., 2023b; Liu et al., 2023).
-Therefore, in MMRec, item IDs are often used as auxiliary information and combined with multimodal data to enhance item representations(Wang et al., 2024; Zhang et al., 2024e).
-This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving the quality of recommendations.
-A typical MMRec training process generally involves feature extraction and modality fusion (He and McAuley, 2016; Wei et al., 2019; Wang et al., 2021; Tao et al., 2022).
-Feature extraction (Guo et al., 2019) aims to obtain representative features from multiple modalities to better capture the attributes and preferences of users and items.
-Modality fusion (Gao et al., 2020) focuses on effectively integrating data from different modalities to enhance the overall performance of the MMRec and the relevance of its results.
-The recommendation models (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023) combine multimodal data with user-item interaction data to predict user preferences and provide personalized recommendations.
-2.3. Foundation Models
-Foundation Models (FM) are often pre-trained on large-scale unlabeled datasets for general purposes, that can be adapted to a variety of downstream tasks through appropriate fine-tuning, such as task-specific training (Bommasani et al., 2021).
-This makes them highly generalizable and adaptable across different applications.
-They are then adapted to specific recommendation tasks through methods such as representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), model fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), and prompt tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024).
-Some studies have explored the use of FM to enhance the content-awareness of MMRec, improving content understanding and generation by leveraging alignment and generation techniques across multiple modalities.
-CLIP (Radford et al., 2021a; Wu et al., 2024) uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain multimodal representations, capturing the matching relationships between users and items.
-M6-Rec (Cui et al., 2022) builds a recommendation system based on large-scale multimodal pre-trained models specifically designed for the e-commerce domain.
-P5 (Geng et al., 2022) transforms the recommendation task into a natural language processing task, generating personalized recommendations through prompts.
-VIP5 (Geng et al., 2023) leverages pre-trained multimodal models to capture complementary information from visual and textual data, providing more personalized and relevant recommendations to users.
-FedPA (Zhang et al., 2024c) and TransFR (Zhang et al., 2024b) integrate FM with FedRec.
-FedPA introduces a personalized adapter to fine-tune the FM, allowing each client to learn individual user preferences in a lightweight manner.
-In contrast, TransFR leverages FMs to generate universal textual representations, replacing traditional ID-based item representations and enhancing the system’s adaptability to new domains and its ability to handle cold-start issues.
+Traditional Recommendation Systems (RecSys) primarily rely on user-item interactions (Yu et al., 2023).
+In contrast, Multimodal Recommendation systems (MMRec) enhance accuracy and user experience by leveraging multiple data modalities like text, images, audio, and video (Zhou et al., 2023b).
+Item IDs, often used as unique identifiers, can capture collaborative signals and structural relationships (Yuan et al., 2023b; Liu et al., 2023).
+In MMRec, these IDs are frequently combined with multimodal data as auxiliary information to create richer item representations (Wang et al., 2024; Zhang et al., 2024e).
+This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving recommendation quality.
+A typical MMRec process involves two main stages: feature extraction and modality fusion (Guo et al., 2019; Gao et al., 2020).
+Feature extraction (Guo et al., 2019) aims to obtain representative features from various modalities.
+Foundation Models (FMs), typically pre-trained on large-scale unlabeled datasets (Bommasani et al., 2021), have become increasingly prominent for this task.
+They can be adapted for recommendation through various techniques, including representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), direct fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), parameter-efficient adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), or prompt-based tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024).
+Several studies leverage FMs to enhance the content understanding capabilities of MMRec, often using cross-modal alignment or generation techniques.
+CLIP (Radford et al., 2021a; Wu et al., 2024), for example, uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain powerful multimodal representations.
+M6-Rec (Cui et al., 2022) is an FM-based system tailored for e-commerce recommendations.
+P5 (Geng et al., 2022) frames recommendation as a natural language task solvable via prompts.
+VIP5 (Geng et al., 2023) utilizes pre-trained multimodal models to capture complementary visual and textual information for personalized recommendations.
+Following feature extraction, modality fusion (Gao et al., 2020) focuses on effectively integrating these features to enhance overall performance.
+Recommendation models then combine these fused multimodal features with interaction data to predict user preferences (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023).
+However, realizing the full potential of MMRec within the constraints of FL, especially regarding personalized and adaptive fusion across heterogeneous clients, necessitates specialized frameworks are needed to provide personalized recommendation with the privacy protection.
 3. Problem Formulation
-Let $\mathcal{U}=\{u\}$ and =$\mathcal{I}=\{i\}$ be the set of users and items correspondingly, we assume that the users’ interaction history data is implicit feedback, represented by $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n}]^{T}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|},$ where $r_{u,i}=1$ if user $u$ has interacted with item $i$, and $r_{u,i}=0$ otherwise.
-For user $u$, we use $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$ to denote their interaction history and $\mathcal{I}_{u}$ to represent the set of items they have interacted with.
-Moreover, we define $\mathcal{O}=\{(u,i)|\text{user }u\text{ has interacted with item }i\}$ to index the items each user has interacted with.
-Each item in $\mathcal{I}$ has a set of modality features $\mathcal{M}$.
-The $m$-th modality feature is represented as $\mathbf{E}_{m}=[\mathbf{e}_{1,m},\mathbf{e}_{2,m},\dots,\mathbf{e}_{|\mathcal{I}|,m}]$.
-To capture user diversity, FedMR also introduces a set of modality fusion strategies $\mathcal{G}=\{g_{1},g_{2},\cdots,g_{|\mathcal{G}|}\}$.
-In this work, we focus on the visual and textual modalities, denoted as =$\mathcal{M}=\{v,c\}$, respectively.
-It is important to note that FedMR is not limited to these two modalities and the existing fusion strategies, it can include additional modalities or modify the fusion strategies as needed.
-3.1. Objective of FedRec
-In classic FedRec with implicit feedback, each client only has the user’s historical interaction data $\mathbf{r}_{u}$.
-The goal is to train a personalized recommendation model for each user while protecting user privacy, providing personalized recommendations by predicting possible scores $\hat{\mathbf{r}}_{u}$ for items the user has not yet interacted with, based on their observed historical data $\mathbf{r}_{u}$.
-To measure the difference between the prediction $\hat{\mathbf{r}}_{u}$ and the truth $\mathbf{r}_{u}$, we have the following:
-(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}).$
-Here, the function $\mathcal{L}_{\text{recon}}$ measures the reconstruction error between $\hat{\mathbf{r}}_{u}$ and $\mathbf{r}_{u}$, and $\alpha_{u}$ represents the loss weight for the client $u$ ($u\in\mathcal{U}$), which is used to balance the client’s contribution to the overall training, and satisfies the condition $\sum_{u\in\mathcal{U}}\alpha_{u}=1$.
-By optimizing Eq.
-(1), FedRec aims to make the predicted item scores $\hat{\mathbf{r}}_{u}$ for each user $u$ as close as possible to their true interactions ${\mathbf{r}}_{u}$.
+Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively.
+User-item interactions are represented by a binary matrix $\mathbf{R}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|}$, where $r_{u,i}=1$ signifies an interaction between user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, and $r_{u,i}=0$ otherwise.
+The interaction history for user $u$ is denoted by the vector $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$, and the set of observed interactions is $\mathcal{O}=\{(u,i)\mid r_{u,i}=1\}$.
+Each item $i\in\mathcal{I}$ is associated with features from a set of modalities $\mathcal{M}$.
+For the $m$-th modality, these features are represented by a matrix $\mathbf{E}_{m}\in\mathbb{R}^{|\mathcal{I}|\times d_{m}}$, where the vector $\mathbf{e}_{i,m}\in\mathbb{R}^{d_{m}}$ is the feature embedding for item $i$ in modality $m$.
+To handle multimodal information effectively, we introduce a set of modality fusion strategies $\mathcal{G}=\{g_{1},\dots,g_{|\mathcal{G}|}\}$ for dynamically combining features.
+We primarily focus on visual $v$ and textual $c$ modalities, i.e., =$\mathcal{M}=\{v,c\}$, although the FedMR framework is extensible to other modalities and fusion mechanisms.
+4. Methodology
+FedMR aims to seamlessly integrate multimodal item features with ID embeddings within a federated environment, enhancing existing ID-based FedRecs, to predict items that users have not yet explored but may find interesting.
+It centralizes multimodal data processing on the server for efficiency and rich item context, then uses dynamic client-side personalization via MFFM to handle heterogeneity.
+4.1. Preliminaries
+FedRec systems aim to collaboratively train recommendation models while preserving user privacy.
+The standard objective is to minimize the aggregated reconstruction loss across clients $u\in\mathcal{U}$:
+(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}),$
+where $\mathcal{L}_{\text{recon}}$ is the loss function, and $\alpha_{u}$ is the weight of client $u$’s contribution.
+Many conventional FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023d) rely solely on a global item ID embedding matrix $\mathbf{D}$ and user-specific parameters $\theta_{u}$ to predict a score $\hat{r}_{ui}$ using a prediction function $f$:
+(2) $\hat{r}_{ui}=f(\mathbf{D}_{i};\theta_{u}).$
+Here, $\mathbf{D}_{i}$ represents the ID embedding for item $i$.
+These methods, so-called ID-based FedRecs, primarily capture collaborative signals and structural relationships (Yuan et al., 2023b), but often overlook rich multimodal item content, which motivating us to propose the FedMR.
+4.2. Server-Side Multimodal Encoding
+To efficiently incorporate multimodal item content while mitigating client resource limitations and communication overhead, FedMR performs feature encoding centrally on the server.
+Leveraging a pre-trained FM $h(\cdot)$, FedMR encodes raw visual features $\mathbf{E}_{v}$ and textual contents $\mathbf{E}_{c}$ into dense semantic embeddings:
+(3) $\mathbf{V}=h(\mathbf{E}_{v}),\quad\mathbf{C}=h(\mathbf{E}_{c}).$
+This server-side encoding yields communication-efficient, high-quality multimodal representations.
+Alongside these content embeddings $\mathbf{V}$ and $\mathbf{C}$, FedMR maintains the standard item ID embedding matrix $\mathbf{D}$ to capture collaborative filtering signals (Liu et al., 2023).
+The three representations $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ then used to form the input for the subsequent fusion module of FedMR.
+4.3. Mixing Feature Fusion Module
+FedMR introduce a novel Mixing Feature Fusion Module (MFFM), designed to dynamically fuse the server-provided global representations with client-specific context. MFFM operates in two stages:
+Server-Side Multi-Strategy Fusion.
+To capture diverse feature interactions, the server employs a predefined set of $|\mathcal{G}|$ distinct fusion strategies $\{g_{j}\}_{j=1}^{|\mathcal{G}|}$.
+Each strategy $g_{j}$, parameterized by $\gamma_{j}$, combines the multimodal embeddings $\mathbf{V}$ and $\mathbf{C}$, and the ID embedding $\mathbf{D}$ to generate a specific fused representation $\mathbf{F}_{j}$:
+(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}).$
+Using multiple strategies produces a diverse set of feature representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$, offering various perspectives on item characteristics.
+These representations are then transmitted to the clients.
+Client-Side Personalized Refinement.
+To address user heterogeneity, each client $u$ employs a personalized refinement mechanism.
+A user-specific router module $\phi_{u}$, parameterized by local parameters $\varphi_{u}$, dynamically computes importance weights $\mathbf{w}_{u}$ for the received server representations based on the client’s interaction history $\mathbf{r}_{u}$:
+(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}),$
+where $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ is typically normalized.
+Each weight $w_{u,j}$ signifies the relevance of the $j$-th feature perspective $\mathbf{F}_{j}$ for user $u$.
+The final personalized item representation $\bar{\mathbf{F}}^{(u)}$ is obtained by adaptively combining the server representations using these learned weights:
+(6) $\bar{\mathbf{F}}^{(u)}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
+This adaptive weighting, learned locally based on $\mathbf{r}_{u}$, allows FedMR to tailor global representations to individual user preferences and local data characteristics, directly mitigating the impact of non-IID data and enhancing personalization with privacy-preserving.
+4.4. Objective of FedMR
+By substituting the representation $\bar{\mathbf{F}}_{i}^{(u)}$ corresponding to item $i$ for the standard ID embedding $\mathbf{D}_{i}$ within the prediction function $f$ in Eq.
+(2), the overall objective in Eq.
+(1) becomes the joint optimization of global and local parameters to obtain the predicted scores $\hat{r}_{u,i}$:
+(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}_{i}^{(u)};\theta_{u}),r_{u,i}).$
+Eq.
+(7) maintains compatibility with existing FedRec frameworks using prediction functions like $f$, allowing them to benefit from richer, personalized, multimodal item representations while preserving privacy.
+This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
 Figure 1.
 The framework of FedMR.
-FedMR utilizes a FM to process multimodal data, including visual and textual information, to enhance personalized recommendations.
-On the server side, FedMR encodes visual and textual data into fixed embeddings.
-On each client, a Mixing Feature Fusion Module dynamically combines these embeddings with ID embeddings, adjusting the fusion weights based on user interactions to generate personalized item embeddings, capturing fine-grained user diversity.
-3.2. ID-based FedRec
-Currently, most FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023c, d; Li et al., 2023a, 2024b) construct a global item embedding $\mathbf{D}$ to capture the items’ features and their relationships with neighboring items, forming $\hat{\mathbf{r}}_{u}$ as follows:
-(2) $\hat{\mathbf{r}}_{u}=f(\mathbf{D};\theta_{u}),$
-where the function $f$, driven by the user-specific parameters $\theta_{u}$, generates personalized predictions $\hat{\mathbf{r}}_{u}$ upon receiving $\mathbf{D}$.
-However, these methods overlook the rich item multimodal information.
-Therefore, we consider such methods as ID-based FedRec in this work.
-4. Methodology
-The aim of our proposed method FedMR is to integrate the multimodal features of items with ID embeddings within a federated environment, and incorporates these features into existing ID-based FedRec methods.
-By doing so, we transform current FedRec methods into FedMMR methods, allowing them to fully utilize the multimodal features of items to predict items that users have not yet viewed but may find interesting, while protecting user privacy.
-4.1. Encoding Multimodal Information with FM
-In RecSys, users typically interact with only a small subset of all available items, which are usually stored on the server.
-Therefore, in multimodal scenarios, we assume the server holds the complete modal information of the item set.
-FedMR begins by introducing a foundation model $h$ on the server side to encode the items’ modal information, producing feature embeddings for each modality:
-(3) $\mathbf{V}=h(\mathbf{E}_{v}),\quad\mathbf{C}=h(\mathbf{E}_{c}).$
-Eq.
-3 has two main advantages: 1) It transforms the visual or textual modality $\mathbf{E}_{m}$ into a format that the recommendation model can process; 2) Compared to transmitting raw modality, the encoded feature embeddings require less communication overhead, thus improving communication efficiency.
-Moreover, FedMR also builds an id embedding $\mathbf{D}$ to capture the structure information of items.
-4.2. Mixing Feature Fusion Module
-Considering the diverse interests and behavior patterns of users, FedMR incorporates a Mixing Feature Fusion Module (MFFM) that adopts multiple modality fusion strategies $\{g_{j}\}$ from $\mathcal{G}$ to fuse the multimodal feature embeddings $\mathbf{V}\text{, }\mathbf{C}$, and the id embedding $\mathbf{D}$ :
-(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}),$
-where $\gamma_{j}$ is parameters of the fusion strategy $g_{j}$, and $\mathbf{F}_{j}$ is the feature embedding obtained after applying the $j$-th fusion strategy.
-To capture user diversity by enabling each strategy to focus on specific user groups based on their interaction histories $\mathbf{r}_{u}$, FedMR introduces a user-specific router $\phi_{u}$, driven by the parameters $\varphi_{u}$:
-(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}).$
-The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$.
-Then, we have the final fused features $\bar{\mathbf{F}}^{(u)}$ as follows:
-(6) $\bar{\mathbf{F}}^{(u)}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
-According to Eq.
-(6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item representations $\bar{\mathbf{F}}^{(u)}$, thereby enhancing the system’s ability to capture fine-grained user preferences.
-4.3. Objective of FedMR
-By replacing $\mathbf{D}$ in Eq. (2) with $\bar{\mathbf{F}}$ obtained from Eq. (6) and then substituting into Eq. (1), we have the objective of FedMR as follows:
-(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}^{(u)};\theta_{u}),r_{u,i}).$
-Eq.
-(7) defines the objective of FedMR and is compatible with most existing ID-based FedRec methods.
-By replacing the original reconstruction error calculation function and the values of $\{\alpha_{u}\}$ in these methods with Eq.
-(7), they can integrate the multimodal information of items into personalized recommendations while protecting user privacy.
-This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
-4.4. Framework
-Fig.
-1 illustrates the overall framework of FedMR, which leverages a FM to process multimodal information for enhancing personalized recommendations.
-On the server side, FedMR uses item cover images as visual modality data and titles as textual modality data.
-These inputs are processed by the FM to generate fixed visual embeddings $\mathbf{V}$ and text embeddings $\mathbf{C}$.
-Since both the FM and the input data are fixed, these item feature embeddings do not need to be updated during training.
-Additionally, the server maintains an ID embedding $\mathbf{D}$ to capture the structural information of the items.
-Upon receiving the features $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM.
-Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion.
-The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy.
-The personalized item representations $\bar{\mathbf{F}}^{(u)}$ are then fed into a prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$, and optimize the model parameters by minimizing reconstruction loss in Eq.
-(7).
-During the entire process, only the ID embedding $\mathbf{D}$ and the gradients of fusion strategies in the MFFM are sent back to the server for aggregation and updates, hereby protecting user privacy while integrating shared information across clients.
-The core of FedMR lies in dynamically adjusting weights in the MFFM based on the user’s local interactions $\mathbf{r}_{u}$, thereby enhancing the diversity and personalization of the RecSys while ensuring privacy protection.
-As a result, FedMR is compatible with existing ID-based FedRec methods, allowing multimodal item information to be integrated into these methods.
-Furthermore, FedMR can be extended by incorporating more modalities or adjusting fusion strategies as needed.
+It comprises two components: (1) Server-Side Multi-Strategy Fusion, which fuses multimodal features on server, and (2) Client-Side Personalized Refinement, which dynamically adjusts representations based on user finteractions for personalized recommendations on each client while preserving user’s data privacy.
 Algorithm 1 FedMR
@@
 Initialize: $\{\theta_{u}\}$, $\{\varphi_{u}\}$, $\{\gamma_{j}\}$, $\mathbf{D}$
-GlobalProcedure:
+MultiStrategyFusion:
 1:$\mathbf{V}\leftarrow h(\mathbf{E}_{v})$, $\mathbf{C}\leftarrow h(\mathbf{E}_{c})$ according to Eq. (3);
 2:for $t=1,2,\dots,T$ do
-3:     $S_{t}$leftarrow randomly select $n_{s}$ from $n$ clients;
-4:     for all client index $u\in S_{t}$ do
-5:        $\hat{\mathbf{r}}_{u}$, $\mathbf{D}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}$leftarrow ClientUpdate($\varphi$, $\theta$);
-6:     end for
-7:     $\mathbf{D}\leftarrow\sum_{u\in S_{t}}\alpha_{u}\mathbf{D}^{(u)}$;
-8:     $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
-9:  end for
-10:  return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
-ClientUpdate:
-1:  $\nabla^{(u)}_{j}{\gamma}\leftarrow 0$, $j=1,2,\cdots|\mathcal{G}|$;
-2:  for $a=1,2,\dots,A$ do
-3:     Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}$ by differentiating Eq. (7);
-4:     Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$;
-5:     Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$;
-6:     Update $\mathbf{D}\leftarrow\mathbf{D}-\eta\nabla_{\mathbf{D}}$;
-7:     Update $\gamma_{j}\leftarrow\gamma_{j}-\eta\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
-8:     Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
+3:    $S_{t}\leftarrow$ randomly select $n_{s}$ from $n$ clients;
+4:    Compute $\mathbf{F}_{j}$ according to Eq. (4), $j=1,2,\cdots|\mathcal{G}|$;
+5:    for all client index $u\in S_{t}$ do;
+6:         $\hat{\mathbf{r}}_{u}$, $\nabla_{\mathbf{D}}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}\leftarrow\text{ClientUpdate}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|})$;
+7:    end for
+8:    $\mathbf{D}\leftarrow\mathbf{D}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla_{\mathbf{D}}^{(u)}$;
+9:    $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
+10:end for
+11:return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
+PersonalizedRefinement:
+1:$\nabla_{\mathbf{D}}^{(u)}\leftarrow 0$; $\triangleright$ Initialize gradient accumulators
+[…]
+3:for $a=1,2,\dots,A$ do
+4:    Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla_{\gamma_{j}}\}$ by differentiating Eq. (7);
+5:    Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$triangleright Update local prediction parameters;
+6:    Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$triangleright Update local router parameters;
+7:    Accumulate $\nabla_{\mathbf{D}}^{(u)}\leftarrow\nabla_{\mathbf{D}}^{(u)}+\nabla_{\mathbf{D}}$triangleright Accumulate gradients;
+8:    Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
 9:    Compute $\hat{\mathbf{r}}_{u}$ according to Eq. (2);
 10:end for
-11:  return: $\hat{\mathbf{r}}_{u}$, $\mathbf{D}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
-4.5. Algorithm
-To optimize the FedMR objective function defined in Eq. (7), we propose an alternating optimization algorithm for training FedMR. The detailed steps of the algorithm are presented in Alg. 1.
-The FedMR framework consists of two main components: the GlobalProcedure and the ClientUpdate.
-In the GlobalProcedure, the server first uses a FM $h(\cdot)$ to encode the multimodal data, generating the corresponding embeddings $\mathbf{V}$ and $\mathbf{C}$.
-The server then randomly selects a subset of clients to participate in each training round to reduce the risk of privacy leakage caused by clients participating in consecutive rounds (Li et al., 2023a).
-For each selected client $u\in S_{t}$, the client update process is executed to obtain the local prediction result $\hat{\mathbf{r}}_{u}$, the local updated id embeddings $\mathbf{D}^{(u)}$, and the local accumulated gradient $\{\nabla^{(u)}_{j}\gamma\}$.
-The server aggregates the results returned by the clients in a weighted manner, and updates the global shared embedding $\mathbf{D}$ and fusion strategy parameters $\{\gamma_{j}\}$.
-During the ClientUpdate, each client calculates the gradients of the model parameters, such as $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$, and $\nabla^{(u)}_{\gamma_{j}}$, based on the FedMR objective function in Eq.
-(7) and updates these parameters using gradient descent (Amari, 1993).
-Additionally, each client accumulates the local gradient $\nabla^{(u)}_{\gamma_{j}}$ into $\nabla^{(u)}_{j}\gamma$ for the global update of the fusion strategy parameters.
-By incorporating a FM $h(\cdot)$ to handle multimodal data $\mathcal{M}$, FedMR achieves efficient personalized multimodal recommendation while preserving user privacy.
+11:return: $\hat{\mathbf{r}}_{u}$, $\nabla_{\mathbf{D}}^{(u)}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
+4.5. Unified Framework and Algorithm
+The FedMR framework, illustrated in Fig.
+1, integrates multimodal information into FedRecs while preserving privacy through an alternating optimization procedure, as shown in Alg.
+1).
+In each communication round $t$, the server first computes the multi-strategy fused representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$ in Eq.
+(4) using current global parameters ($\mathbf{D},\{\gamma_{j}\}$) and transmits them to a randomly selected client subset $S_{t}$.
+Each client $u\in S_{t}$ then utilizes its personalized router $\phi_{u}$ in Eq.
+(5) to compute user-specific weights $\mathbf{w}_{u}$ based on its local history $\mathbf{r}_{u}$, generating the personalized item representation $\bar{\mathbf{F}}^{(u)}$ in Eq.
+(6).
+Using $\bar{\mathbf{F}}^{(u)}$, the client performs local updates on its parameters $\theta_{u}$ and $\varphi_{u}$ by minimizing the objective in Eq.
+(7) on its private data.
+Concurrently, it computes gradients w.r.t.
+the global parameters, $\nabla_{\mathbf{D}}^{(u)}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}_{j=1}^{|\mathcal{G}|}$.
+These gradients are accumulated and sent back to the server.
+Finally, the server aggregates the received gradients from all clients in $S_{t}$, typically weighted by $\alpha_{u}$, and updates the global parameters $\mathbf{D}$ and $\{\gamma_{j}\}$.
+This iterative exchange enables collaborative learning.
+Crucially, the client-side personalized refinement directly addresses data heterogeneity, while privacy is maintained as only gradients are communicated.
+FedMR’s design thus flexibly integrates multimodal data, enables fine-grained personalization, maintains compatibility with ID-based methods, and ensures efficiency and accuracy within privacy constraints.
+4.6. Convergence Analysis
+Here, we provide a theoretical analysis of the FedMR under assumptions for non-convex federated optimization.
+For simplicity, we denote our learnable parameters as $\mathbf{w}=(\Theta_{g},\{\Theta_{l,u}\}_{u\in\mathcal{U}})$, where $\Theta_{g}=\{\mathbf{D},\{\gamma_{j}\}\}$ represents the set of global parameters, and $\Theta_{l,u}=\{\theta_{u},\varphi_{u}\}$ denotes the local parameters for client $u$, capturing user-specific prediction and router parameters.
+The global objective is:
+(8) $F(\mathbf{w})\;=\;\sum_{u\in\mathcal{U}}\alpha_{u}\,F_{u}(\mathbf{w}),$
+where each $F_{u}(\mathbf{w})$ is the local objective associated with user $u$, and $\alpha_{u}>0$ with $\sum\alpha_{u}=1$. We aim to find $\mathbf{w}$ that minimizes $F(\mathbf{w})$.
+To establish the convergence guarantees, we rely on the following assumptions commonly used in the analysis of FL algorithms:
+Assumption 1 (Smoothness).
+Each local objective $F_{u}(\mathbf{w})$ is $L$-smooth w.r.t. $\mathbf{w}$. Hence, the global objective $F$ is also $L$-smooth:
+(9) $\|\nabla F_{u}(\mathbf{w})-\nabla F_{u}(\mathbf{w}^{\prime})\|\;\leq\;L\,\|\mathbf{w}-\mathbf{w}^{\prime}\|.$
+Assumption 2 (Bounded Gradient Variance).
+For the stochastic gradient $g_{l,u}(\mathbf{w};\xi)$ of $F_{u}$, we have $\mathbb{E}_{\xi}[g_{l,u}]=\nabla_{\Theta_{l,u}}F_{u}(\mathbf{w})$ and $\mathbb{E}_{\xi}[\|g_{l,u}-\nabla_{\Theta_{l,u}}F_{u}(\mathbf{w})\|^{2}]\leq\sigma_{l}^{2}$.
+Furthermore, there exists $G>0$ such that $\|\nabla F_{u}(\mathbf{w})\|\leq G$ in expectation across clients.
+Assumption 3 (Bounded Heterogeneity).
+The variance of all local gradients across clients is bounded as follows:
+$\mathbb{E}_{u}\bigl{[}\|\nabla F_{u}(\mathbf{w})-\nabla F(\mathbf{w})\|^{2}\bigr{]}\;\leq\;\sigma^{2}.$
+These assumptions ensure that each $F_{u}$ is reasonably well-behaved and client heterogeneity is bounded. Next, we formalize local drift lemmas and provide the main theorem on FedMR’s convergence.
+Let $A$ be the local number of SGD steps per communication round.
+During each round $t$, a subset of clients $\{u\in S_{t}\}$ updates their local parameters $\Theta_{l,u}$ using a local learning rate $\eta$, while the global parameters $\Theta_{g}$ remain fixed until aggregation.
+We can show:
+Lemma 0 (Local Drift Bound).
+Under Assumptions 1 and 2, the expected deviation in local parameters after $A$ updates is bounded by a term proportional to $A\eta^{2}(G^{2}+\sigma_{l}^{2})$.
+Lemma 1 induces a gradient bias when uploading local updates to the server. Combining the drift bound with smoothness yields:
+Lemma 0 (Gradient Bias Bound).
+Let $\bar{g}_{g}^{t}$ denote the averaged gradient for global parameters $\Theta_{g}$ after $A$ local updates on the sampled clients, and $B_{t}=\mathbb{E}[\bar{g}_{g}^{t}|\mathbf{w}^{t}]-\nabla_{\Theta_{g}}F(\mathbf{w}^{t})$ be the gradient bias due to local updates relative to the global parameters $\mathbf{w}^{t}$ at the start of the round.
+Under Assumption 1 and Lemma 1, We then have:
+(10) $\mathbb{E}[\|B_{t}\|^{2}]\leq L^{2}C_{1}A\eta^{2}(G^{2}+\sigma_{l}^{2}).$
+We now can show FedMR converges to a stationary point of $F(\mathbf{w})$ in a manner consistent with typical non-convex FL rates:
+Theorem 3 (Convergence of FedMR).
+Suppose Assumptions 13 hold and let $\eta=c_{\eta}/\sqrt{T}$ for a suitable constant $c_{\eta}>0$. Then for $T$ communication rounds, the FedMR algorithm satisfies:
+(11) $\small\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\bigl{[}\|\nabla F(\mathbf{w}^{t})\|^{2}\bigr{]}\;\leq\;O\Bigl{(}\frac{1}{\sqrt{T}}\Bigr{)}\;+\;O\Bigl{(}\frac{AG^{2}}{T}\Bigr{)}\;+\;O\Bigl{(}\frac{\sigma^{2}+G^{2}}{\sqrt{T}\,n_{s}}\Bigr{)}.$
+Hence, FedMR achieves a convergence rate of $O(T^{-1/2})$ to a stationary point under non-convex settings, matching the standard rates in federated optimization.
+The above theorem indicates that FedMR’s multimodal feature fusion and personalization do not degrade the asymptotic rate, and it remains comparable to baseline FL methods.
+The result also quantifies how larger local steps $A$, stronger data heterogeneity $\sigma^{2}$, or fewer sampled clients $n_{s}$ can slow practical convergence.
+Nonetheless, for moderate $A$, stable $\sigma^{2}$, and typical $n_{s}$, FedMR effectively converges and benefits from the rich multimodal item representations under the federated setup.
 4.7. Complexity Analysis
-Since FedMR is compatible with various FedRec methods, we focus on the time and space complexity introduced by incorporating multimodal data of items in this work.
-Assuming $n$ users, $m$ items and $q$ experts, FedMR , where the multimodal information of the items is encoded on the server side into feature embeddings $\mathbf{V}\in\mathbb{R}^{m\times d}$ and $\mathbf{C}\in\mathbb{R}^{m\times d}$ using a foundation model, with $d\ll\{n,m\}$.
-Since these embeddings remain fixed during training, they contribute a space complexity of $O(md)$ while having negligible impact on time complexity.
-FedMR also maintains an ID embedding $\mathbf{D}\in\mathbb{R}^{m\times d}$ to capture structural information of the items, with a space complexity of $O(md)$.
-During each communication round, the server aggregates and updates local data from the clients, requiring a time complexity of $O(\alpha_{u}md(q+1))$, while each client performs the Mixing Feature Fusion Module to fuse features, resulting in a total client-side space complexity of $O(md+K)$ and a time complexity of $O(P)$, where $K$ and $P$ represents the size and cost time of the component parameters, such as Router, MLP, Gate, etc, respectively.
-Therefore, FedMR requires a total space complexity of $O(n(md+K)+3md)$ and a total time complexity of $O(nmd(q+1)+nP)$.
+In FedMR, the server initially encodes multimodal features into fixed embeddings $\mathbf{V},\mathbf{C}\in\mathbb{R}^{|\mathcal{I}|\times d}$ using an FM $h(\cdot)$, contributing primarily $O(|\mathcal{I}|d)$ space complexity each, as this is often pre-computed.
+The server also stores the learnable ID embedding $\mathbf{D}\in\mathbb{R}^{|\mathcal{I}|\times d}$, adding $O(|\mathcal{I}|d)$ space.
+In each round, the server computes $|\mathcal{G}|$ fused representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$ and aggregates gradients, and the dominant per-round time complexity on the server is $O(|\mathcal{G}||\mathcal{I}|d)$.
+On the client side, storing local parameters $K$ and temporarily holding the $|\mathcal{G}|$ received representations requires approximately $O(|\mathcal{I}|d+K)$ space.
+The client’s per-iteration time complexity for local updates, involving the prediction function $f$ and the MFFM router $\phi_{u}$, is denoted as $O(P)$.
+Therefore, the total system space complexity is roughly $O(|\mathcal{U}|(|\mathcal{I}|d+K)+|\mathcal{I}|d)$, and the total time complexity per round involves server computation plus aggregated client computation, scaling approximately as $O(|\mathcal{G}||\mathcal{I}|d+n_{s}AP)$ where $n_{s}=|S_{t}|$ and $A$ is the number of local steps.
+This analysis indicates manageable overhead for incorporating FedMR to achieve personalized and privacy-preserving multimodal recommendation.
+4.8. Privacy Preservation
+FedMR ensures user privacy primarily through the inherent data localization of the federated learning paradigm, where raw interaction data $\mathbf{r}_{u}$ remains on the client’s device.
+Communication is limited to model gradients or updates.
+The MFFM is designed to preserve this standard: server-side computations involve only global representations, while the client-specific router $\phi_{u}$ and its parameters $\varphi_{u}$ remain entirely local.
+The gradients $\nabla\mathbf{D}^{(u)}$ and $\{\nabla^{(u)}_{j}\gamma\}$ shared with the server are analogous to those shared in standard FedRecs, introducing no additional direct data leakage channels and aligning with current privacy practices (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024c; Feng et al., 2024; Wu et al., 2024).
+Privacy is further enhanced by standard techniques like random client sampling in each round (Chai et al., 2020; Li et al., 2023a).
+Moreover, FedMR is fully compatible with advanced privacy-enhancing technologies such as differential privacy (DP) (Abadi et al., 2016), which can be applied by adding noise to gradients before transmission, and other secure aggregation protocols.
+Our ablation studies provide empirical evidence that DP can be incorporated without drastically compromising the performance.
+5. Experiments
+5.1. Datasets
+We evaluate FedMR’s performance on four public datasets from the NineRec benchmark [^1] Publicly available at: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a): KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie).
+Detailed statistics are provided in Table 1.
+The KU dataset contains short video interactions from the KuaiShou platform.
+The Food, Dance, and Movie datasets comprise video interactions from distinct categories on the BiliBili platform.
+These datasets feature diverse content and user interaction patterns but share the challenge of extreme data sparsity (over 95%).
+We chose these four datasets from the NineRec benchmark for comprehensive evaluation.
+They represent diverse real-world scenarios, originating from two distinct platforms and covering varied content domains at different scales.
+This diversity, coupled with their inherent high sparsity, provides a robust testbed for evaluating FedMR’s adaptability and performance across challenging, realistic conditions.
+Following standard practice, we filter out users with fewer than 5 interactions to mitigate cold-start issues.
+For data preprocessing, missing item images are imputed using the average visual features of existing items, and missing item titles are replaced with the placeholder text ”The title is missing.”
 Table 1.
@@
 #Items indicates the number of items.
+Sparsity is the percentage of #Interactions out of the total possible ratings.
 Dataset #Users #Items #Interactions Sparsity
@@
 Movie 15908 2528 111091 99.72%
-4.7. Privacy Preservation
-FedMR uses a distributed architecture based on FL to effectively reduce the risk of user privacy leakage.
-Except for the Mixing Feature Fusion Module, all other components of FedMR have the same privacy risk level as existing ID-based federated recommendation methods.
-In the Mixing Feature Fusion Module, only the locally updated ID embeddings $\mathbf{D}^{(u)}$ and the local accumulated gradients ${\nabla^{(u)}_{j}\gamma}$ are sent to the server, without revealing any raw user data, also maintaining the same communication privacy protection strategy as current methods (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024b; Feng et al., 2024; Wu et al., 2024).
-To protect user privacy, randomly selecting participants in each communication round and avoiding the same participants in consecutive rounds can effectively reduce the risk of privacy leakage (Chai et al., 2020; Li et al., 2023a).
-Additionally, FedMR can incorporate techniques such as differential privacy (Abadi et al., 2016) and stochastic gradient clipping to further enhance privacy protection.
-We also conduct an ablation study to evaluate the effectiveness of FedMR when additional privacy-preserving measures are applied.
-5. Experiments
-5.1. Datasets
-In this study, we conducte comprehensive experiments on four downstream datasets from Ninerec [^1] The dataset publicly available on: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a) to evaluate the performance and effectiveness of the proposed method FedMR: KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie).
-Table 1 provides the detailed statistics of these datasets.
-The KU dataset is derived from KuaiShou, a well-known short video platform in China, where each entry represents a short video.
-The Food, Dance, and Movie datasets are collected from the BiliBili platform, focusing on food and cooking videos, dance-related videos, and movie-related videos, respectively.
-Each dataset presents unique content types and user interaction patterns, but all share a common challenge of high sparsity, with sparsity levels exceeding 95%.
-For our experiments, we include only users who had interacted with at least 5 items in each dataset to avoid the cold-start problem.
-Additionally, for missing images in the datasets, we fill them with the average values of visual features; for missing titles, we set ”The title is missing.”
+5.2. Baselines
+To evaluate the performance of FedMR comprehensively, we compare it with three state-of-the-art centralized MMRecs and five widely used ID-based FedRecs for recommendation tasks with implicit feedback: (1) VBPR (He and McAuley, 2016): Incorporates visual features into personalized ranking to enhance recommendation accuracy for implicit feedback.
+(2) BM3 (Zhou et al., 2023a): Integrates multimodal features using a self-supervised learning framework.
+(3) MGCN (Yu et al., 2023): Employs a multi-view graph convolutional network to capture multimodal features of users and items.
+(4) FedAvg (McMahan et al., 2017): The foundational federated learning method, using weighted aggregation based on local client data sizes.
+(5) FCF (Ammad-Ud-Din et al., 2019): Combines collaborative filtering with federated learning, performing equal-weighted average aggregation across clients.
+(6) FedNCF (Perifanis and Efraimidis, 2022): Updates user embeddings locally on clients and synchronizes item embeddings globally on the server.
+(7) PFedRec (Zhang et al., 2023c): Aggregates client-side item embeddings on the server to enable bilateral personalized recommendations.
+(8) FedRAP (Li et al., 2023a): Achieves fine-grained personalization for users and items through additive modeling.
+The centralized MMRec methods VBPR, BM3 and MGCN serve as benchmarks to gauge the performance potential achievable without federated constraints, highlighting the gap federated approaches aim to close.
+The selected four federated methods represent common ID-based FedRec approaches.
+We compare the vanilla versions of the ID-based FedRec baselines against the versions enhanced with FedMR (w/ FedMR) to directly measure the performance gains provided by our framework’s multimodal capabilities [^2] We do not include direct comparisons with AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a) as their official code implementations were not publicly available at the time of our experiments, preventing a fair and reproducible comparison..
+5.3. Experimental Setting
+Following prior work (Zhang et al., 2023c; Li et al., 2023a), we employ negative sampling during training and use the leave-one-out strategy for validation and testing.
+Hyperparameters for all baselines are tuned via grid search on the validation set, including the learning rate $\eta$ within $\{10^{i}\mid i=-4,\ldots,-1\}$ and method-specific parameters.
+We use CLIP (Radford et al., 2021a) as the common foundation model for extracting visual and textual features, implemented via OpenCLIP (Ilharco et al., 2021) with pre-trained ViT-B-32 weights from OpenAI (Radford et al., 2021b; Schuhmann et al., 2022).
+A linear mapping layer projects multimodal and ID embeddings to a shared latent dimension of 64 for efficiency.
+The training batch size is 2048.
+All models incorporating hidden layers use a three-layer MLP structure.
+For federated methods, clients perform $5$ local training epochs per communication round using a consistent aggregation strategy for both vanilla and FedMR-integrated versions.
+5.4. Evaluation
+In our experiments, we treat the observed interactions as positive samples and the rest as negative samples.
+We sort the predicted scores between users and all candidate items in descending order, masking the observed positive samples from the training set.
+Then, we evaluate the recommendation performance using two widely-used metrics for implicit feedbacks: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020).
+In the experiments, we set $K=50$.
 Table 2.
@@
 vanilla 0.1127 0.0562 0.0782 0.0201 0.0696 0.0191 0.0552 0.0158
-FedAvg w/ FedMR 0.5147 (356.70% $\uparrow$) 0.1491 (165.30% $\uparrow$) 0.1007 (28.77% $\uparrow$) 0.0261 (29.85% $\uparrow$) 0.0948 (36.21% $\uparrow$) 0.0268 (40.31% $\uparrow$) 0.0564 (2.17% $\uparrow$) 0.0160 (1.27% $\uparrow$)
+FedAvg w/ FedMR 0.5147 0.1491 0.1007 0.0261 0.0948 0.0268 0.0564 0.0160
+Impr. 356.70% $\uparrow$ 165.30% $\uparrow$ 28.77% $\uparrow$ 29.85% $\uparrow$ 36.21% $\uparrow$ 40.31% $\uparrow$ 2.17% $\uparrow$ 1.27% $\uparrow$
 vanilla 0.2255 0.0643 0.0676 0.0177 0.0820 0.0219 0.0581 0.0163
-FCF w/ FedMR 0.5539 (145.63% $\uparrow$) 0.1597 (148.37% $\uparrow$) 0.0699 (3.40% $\uparrow$) 0.0182 (2.75% $\uparrow$) 0.0918 (11.95% $\uparrow$) 0.0247 (12.79% $\uparrow$) 0.0667 (14.80% $\uparrow$) 0.0180 (10.44% $\uparrow$)
+FCF w/ FedMR 0.5539 0.1597 0.0699 0.0182 0.0918 0.0247 0.0667 0.0180
+Impr. 145.63% $\uparrow$ 148.37% $\uparrow$ 3.40% $\uparrow$ 2.75% $\uparrow$ 11.95% $\uparrow$ 12.79% $\uparrow$ 14.80% $\uparrow$ 10.44% $\uparrow$
 vanilla 0.1176 0.0249 0.0221 0.0055 0.0548 0.0163 0.0185 0.0054
-FedNCF w/ FedMR 0.1961 (66.75% $\uparrow$) 0.0609 (144.58% $\uparrow$) 0.0708 (220.36% $\uparrow$) 0.0187 (240.00% $\uparrow$) 0.0843 (53.83% $\uparrow$) 0.0243 (49.08% $\uparrow$) 0.0406 (119.46% $\uparrow$) 0.0125 (131.48% $\uparrow$)
+FedNCF w/ FedMR 0.1961 0.0609 0.0708 0.0187 0.0843 0.0243 0.0406 0.0125
+Impr. 66.75% $\uparrow$ 144.58% $\uparrow$ 220.36% $\uparrow$ 240.00% $\uparrow$ 53.83% $\uparrow$ 49.08% $\uparrow$ 119.46% $\uparrow$ 131.48% $\uparrow$
 vanilla 0.0637 0.0230 0.0252 0.0068 0.0327 0.0093 0.0170 0.0042
-PFedRec w/ FedMR 0.2157 (238.62% $\uparrow$) 0.0616 (167.83% $\uparrow$) 0.0392 (55.56% $\uparrow$) 0.0092 (35.29% $\uparrow$) 0.0836 (155.66% $\uparrow$) 0.0207 (122.58% $\uparrow$) 0.0413 (142.94% $\uparrow$) 0.0132 (214.29% $\uparrow$)
+PFedRec w/ FedMR 0.2157 0.0616 0.0392 0.0092 0.0836 0.0207 0.0413 0.0132
+Impr. 238.62% $\uparrow$ 167.83% $\uparrow$ 55.56% $\uparrow$ 35.29% $\uparrow$ 155.66% $\uparrow$ 122.58% $\uparrow$ 142.94% $\uparrow$ 214.29% $\uparrow$
 vanilla 0.4706 0.1086 0.0682 0.0186 0.0696 0.0191 0.0453 0.0136
-Fed FedRAP w/ FedMR 0.4824 (2.51% $\uparrow$) 0.1305 (20.17% $\uparrow$) 0.0895 (31.19% $\uparrow$) 0.0247 (32.91% $\uparrow$) 0.1142 (64.08% $\uparrow$) 0.0294 (53.93% $\uparrow$) 0.0586 (29.36% $\uparrow$) 0.0199 (46.32% $\uparrow$)
-5.2. Baselines
-To comprehensively evaluate the performance of FedMR, we select 3 cutting-edge centralized MMRec methods, and 5 ID-based federated methods commonly used in FedRec tasks with implicit feedback to serve as baselines in our study:
-VBPR (He and McAuley, 2016) incorporates visual features into personalized ranking to improve recommendation accuracy based on implicit feedback.
-BM3 (Zhou et al., 2023a) effectively integrates multimodal features through a self-supervised learning framework.
-MGCN (Yu et al., 2023) utilizes a multi-view graph convolutional network to separately capture the multimodal features of users and items.
-FedAvg (McMahan et al., 2017) is the earliest federated learning method, performing weighted aggregation based on the relative sizes of local client data.
-FCF (Ammad-Ud-Din et al., 2019) introduces collaborative filtering into federated learning, assigning equal weights to all clients for average aggregation.
-FedNCF (Perifanis and Efraimidis, 2022) updates user embeddings locally on each client and synchronizes item embeddings on the server for global updates.
-PFedRec (Zhang et al., 2023c) aggregates item embeddings from clients on the server side and uses them during evaluation to achieve bilateral personalized recommendations.
-FedRAP (Li et al., 2023a) achieves fine-grained personalized modeling of users and items through additive personalization.
-The methods VBPR, BM3, and MGCN are three centralized benchmark MMRec methods that demonstrate baseline performances of the used datasets in a centralized recommendation setting.
-The methods FedAvg, FCF, FedNCF, PFedRec, and FedRAP are five ID-based FedRec methods.
-We conducted experiments on both the vanilla versions of these ID-based FedRec methods and their multimodal versions integrated with our proposed method FedMR to demonstrate the effectiveness of FedMR.
-Implementations of all used baseline methods are publicly available through their papers.
-All methods are implemented using PyTorch (Paszke et al., 2019), and all experiments are run on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, 3 Quadro RTX 6000 GPUs, and 395 GB of RAM.
-5.3. Experimental Setting
-Following previous studies (Zhang et al., 2023c; Li et al., 2023a), we perform negative sampling for each positive sample in the training set and use the leave-one-out strategy for validation.
-We conduct hyperparameter tuning for all baseline methods on each dataset used, selecting the learning rate $\eta$ and adjusting the hyperparameters in FedRAP’s objective function from $\{10^{i}\mid i=-4,\ldots,-1\}$.
-To ensure fairness, we use the CLIP (Radford et al., 2021a) implemented by the Open CLIP (Ilharco et al., 2021; Radford et al., 2021b; Schuhmann et al., 2022) as the foundation model in our experiments, loading the ViT-B-32 parameters trained by OpenAI to process multimodal information.
-It should be noted that the foundation model in FedMR can be replaced with other models as needed.
-Additionally, we add a feature mapping layer before the input fusion layer of the multimodal embeddings and ID embeddings, mapping the feature dimensions output by the foundation model to a fixed latent embedding dimension, i.e.
-64, to improve storage efficiency.
-We fix the training batch size at 2048 for all methods, and for methods with hidden layers, the number of layers is set to 3.
-For each federated baseline, we adopt the same aggregation strategy to update the global part of FedMR.
-In the main experiments, all clients participated in the training of both the original versions and the FedMR versions of all federated baselines, and no additional privacy protection measures are applied.
-Additionally, for these federated baselines, the number of local training epochs was set to 10.
-Except for loading the foundation model, we do not use any pre-training strategies for any methods.
-5.4. Evaluation
-In our experiments, we treat the observed interactions as positive samples and the rest as negative samples.
-We sort the predicted interaction scores between users and all candidate items in descending order, masking the observed positive samples from the training set.
-Then, we evaluate the recommendation performance using two widely-used metrics: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020).
-In the experiments, we set $K=50$.
+Fed FedRAP w/ FedMR 0.4824 0.1305 0.0895 0.0247 0.1142 0.0294 0.0586 0.0199
+Impr. 2.51% $\uparrow$ 20.17% $\uparrow$ 31.19% $\uparrow$ 32.91% $\uparrow$ 64.08% $\uparrow$ 53.93% $\uparrow$ 29.36% $\uparrow$ 46.32% $\uparrow$
 5.5. Comparison Analysis
-Table 2 shows the performance of both centralized and federated methods across the four datasets.
-The results highlight that FedMR significantly boosts the performance of the used FedRec methods, particularly in the KU dataset, where HR@50 for FedAvg, FCF, PFedRec, and FedRAP exhibits notable improvements.
-This indicates that FedMR effectively enhances traditional ID-based FedRec methods by integrating multimodal item information.
-The performance improvements in other datasets further demonstrate FedMR’s ability to alleviate data sparsity by better leveraging users’ interaction histories alongside multimodal data for more personalized recommendations, while protecting user privacy.
-Compared to centralized MMRec methods, FedRec approaches using FedMR achieve similar or even superior performance while maintaining privacy, e.g., in the KU dataset, FedAvg with FedMR surpasses all centralized methods in terms of HR@50.
-Federated methods with FedMR also perform strongly across other datasets, reinforcing FedMR’s capability to improve recommendation quality without sacrificing user privacy.
-Table 2 further shows that by incorporating multimodal information, FedMR provides a more comprehensive representation of item features.
-Compared to ID-based FedRec, the use of multimodal data enables more accurate predictions of user preferences.
-Additionally, the MFFM of FedMR dynamically adjusts the weights of visual, textual, and ID features based on each user’s dependence on these modalities.
-This flexible approach to feature fusion avoids the drawbacks of single fusion strategy methods, allowing the model to better capture complex relationships between modalities and further enhance personalized recommendation accuracy.
+The performance comparison detailed in Table 2 reveals key insights into FedMR’s effectiveness relative to both ID-based federated baselines and centralized multimodal systems.
+The results consistently demonstrate FedMR’s significant value in bridging the gap between privacy-preserving FL and effective multimodal recommendation.
+Enhancing ID-based FedRec.
+FedMR consistently and substantially elevates the performance of all evaluated ID-based FedRecs across the four datasets.
+The degree of improvement varies depending on the baseline architecture.
+Performance gains are particularly dramatic for methods initially exhibiting lower performance, such as FedAvg, FedNCF, and PFedRec, especially on datasets where multimodal signals prove highly beneficial like KU and Food.
+This trend suggests FedMR effectively integrates crucial multimodal context and personalized fusion capabilities that compensate for the limitations of simpler ID-based approaches in content understanding and heterogeneity handling.
+While FedMR also enhances more sophisticated baselines like FedRAP, the relative improvement tends to be more moderate, indicating diminishing returns when the base model already incorporates strong personalization mechanisms, though the added multimodal dimension still provides value.
+Performance Relative to Centralized Methods.
+When compared against centralized MMRec benchmarks, FedMR-enhanced methods demonstrate encouraging results, significantly narrowing the typical performance gap observed between federated and centralized systems.
+Notably, on the KU dataset, FedMR integrated with FedAvg and FCF achieves performance surpassing the centralized counterparts.
+This suggests that FedMR’s adaptive client-side fusion can be highly effective, potentially adapting better to user nuances than global centralized fusion in certain scenarios.
+However, a performance gap often persists on the larger Dance, Food, and Movie datasets compared to the best centralized methods.
+This trend indicates that while FedMR offers substantial improvements, the inherent challenges under the federated recommmendation settings, such as data fragmentation and privacy constraints limiting information sharing, still make fully matching centralized performance difficult, particularly on larger-scale, complex datasets.
+Furthermore, the lower absolute scores achieved by all methods on datasets like Dance and Movie suggest these represent intrinsically more challenging recommendation tasks, possibly due to complex user behavior or less discriminative multimodal features.
+Summary.
+In summary, FedMR’s integration of multimodal information through the server-client architecture and the adaptive MFFM consistently improves upon diverse ID-based FedRecs.
+FedMR significantly narrows, and sometimes closes or even surpasses, the performance gap relative to centralized MMRecs.
+The remaining gap on more challenging datasets highlights the ongoing need for advancements in FedMMRec.
+FedMR represents a crucial step in this direction, showcasing the effectiveness of adaptive, personalized fusion mechanisms in unlocking multimodal data’s potential under privacy-preserving federated constraints.
+Specifically, its ability to dynamically adapt to user preferences underscores its value for developing robust and personalized FedMMRecs.
 5.6. Ablation Study
-To validate the effectiveness of each component in FedMR, we conducted experiments using several variants of the system.
-Specifically, we independently applied the Sum, MLP, and Gate strategies for the fusion of multimodal features and ID features, to assess the impact of dynamic weighted fusion based on users’ interaction history $\mathbf{r}_{u}$, $\forall u\in\mathcal{U}$.
-Additionally, we introduce random Gaussian noise into the gradients uploaded to the server to evaluate the potential performance impact of this privacy-preserving measure.
-We also examine the distribution of multimodal and ID features within the same feature space.
-These ablation studies demonstrate the effectiveness of each component of FedMR.
-Figure 2. Visualization of multimodal features and ID features mapped into a shared representation space.
-5.6.1. Ablation Study on Modality Distributions
-To investigate the distribution relationship between multimodal features and ID features, we select FedRAP (Li et al., 2023a) as the backbone for FedMR in this study.
-FedRAP maintains both globally shared and user-specific ID features, which helps demonstrate how different ID features affect the distribution within the feature space.
+To comprehensively evaluate FedMR’s contributions and key design choices, we conduct ablation studies analyzing four crucial aspects.
+We investigate the effectiveness of the MFFM compared to simpler alternatives, assess the individual contributions of visual, textual, and ID modalities, examine the impact of client heterogeneity related to user activity levels, and evaluate the model’s robustness when incorporating noise-based privacy enhancement.
+These analyses collectively highlight the significance of FedMR’s adaptive multimodal fusion and illustrate the interplay between modality utilization, user diversity, and privacy considerations in FedRec.
+Figure 2. Performance comparison of different feature fusion strategies in various FedRecs combined with FedMR on KU.
+5.6.1. Ablation Study on Fusion Strategy
 Fig.
-2 visualizes the distribution of modality features and ID features after training FedRAP with FedMR, on the KU dataset.
-The first subfigure in Fig.
-2 shows that when the multimodal features of items and globally shared ID features are mapped into a shared representation space, they remain in distinct subregions.
-This phenomenon, known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a), helps the model distinguish between different modalities, reducing noise interference and improving the robustness of multimodal data processing.
-The following two subfigures incorporate the personalized ID embeddings of the users 00 and $42$, respectively, highlighting the significant impact of personalized ID features on the distribution in the shared space.
-By comparing the first subfigure with the subsequent two, we observe that the introduction of personalized ID features causes the ID features to gradually expand towards the regions of visual and textual features, shrinking the modality gap.
-This facilitates the fusion of multimodal information and enhances the accuracy of personalized recommendations.
-Furthermore, by comparing the differences in modality gap changes for each user, it becomes evident that the model can adjust the feature distribution based on user-specific behaviors and preferences, thus better satisfying individual user needs.
-This study underscores the importance of personalized ID features in federated recommendation systems.
-Due to privacy constraints, the model cannot access raw user data directly, but through the use of personalized ID features, the model can provide personalized recommendation services while maintaining user privacy.
-Table 3. Performance comparison of FedRec Methods on HR@50 and NDCG@50 with and without multimodal and ID features on the KU dataset.
-Method Metric FedMR w/o $\mathbf{V}$ w/o $\mathbf{C}$ w/o $\mathbf{D}$
+2 shows the HR@50 and NDCG@50 results for five FedRec methods with different fusion strategies on the KU dataset.
+The results highlight the critical role of fusion strategies in performance.
+Among them, the Mix strategy, which dynamically combines fusion methods based on user interaction history, achieves the best performance.
+In contrast, the Sum strategy performs the worst across all metrics.
+The Sum strategy, which simply adds features, fails to account for complex interactions among modalities, often leading to information loss and poor handling of long-tail distributions.
+In contrast, the MLP strategy better captures nonlinear interactions but is limited by its network depth and parameter count.
+The Gate strategy dynamically learns feature weights, optimizing the importance of different modalities.
+However, the Mix strategy combines all fusion methods, dynamically adjusting their contributions, enabling it to generate fine-grained personalized item representations and better capture complex relationships among features.
+This study underscores the importance of personalized item representations in FedMMRec.
+By leveraging the MFFM, FedMR allows each client to dynamically adjust fusion weights based on user interaction history, enabling fine-grained personalization and enhancing recommendation accuracy while maintaining privacy under the federated settings.
+Table 3. Performance comparison of FedRecs with and without multimodal and ID features on KU.
+Method Metric FedMR w/o V w/o C w/o D
 FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.5094 (1.03% $\downarrow$)
 NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1407 (5.66% $\downarrow$)
-FCF HR@50 0.5539 0.1569 (71.67% $\downarrow$) 0.1569 (71.67% $\downarrow$) 0.1275 (76.99% $\downarrow$)
-NDCG@50 0.1597 0.0486 (69.57% $\downarrow$) 0.0476 (70.19% $\downarrow$) 0.0287 (82.04% $\downarrow$)
+FCF HR@50 0.5539 0.2843 (48.67% $\downarrow$) 0.4216 (23.89% $\downarrow$) 0.2353 (57.52% $\downarrow$)
+NDCG@50 0.1597 0.0748 (53.16% $\downarrow$) 0.0965 (39.57% $\downarrow$) 0.0578 (63.81% $\downarrow$)
 FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1602 (18.31% $\downarrow$)
 NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0538 (11.66% $\downarrow$)
-PFedRec HR@50 0.2157 0.2010 (6.82% $\downarrow$) 0.1225 (43.21% $\downarrow$) 0.0490 (77.27% $\downarrow$)
-NDCG@50 0.0616 0.0557 (9.58% $\downarrow$) 0.0341 (44.64% $\downarrow$) 0.0191 (68.97% $\downarrow$)
-FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2647 (30.78% $\downarrow$) 0.1863 (51.29% $\downarrow$)
-NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0734 (43.75% $\downarrow$) 0.0490 (62.43% $\downarrow$)
+PFedRec HR@50 0.2157 0.1990 (7.74% $\downarrow$) 0.1941 (10.01% $\downarrow$) 0.1794 (16.83% $\downarrow$)
+NDCG@50 0.0616 0.0469 (23.86% $\downarrow$) 0.0537 (12.82% $\downarrow$) 0.0473 (23.21% $\downarrow$)
+FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.3235 (15.40% $\downarrow$) 0.1863 (51.29% $\downarrow$)
+NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0714 (45.29% $\downarrow$) 0.0490 (62.43% $\downarrow$)
 5.6.2. Ablation Study on Modality Efficiency
-Table 3 presents a performance comparison of the five used ID-based FedRec methods with the inclusion of multimodal information.
-It evaluates these methods under the full model (FedMR) and with the removal of visual modality (w/o V), textual modality (w/o C), and ID features (w/o D), focusing on HR@50 and NDCG@50 metrics.
-Firstly, removing the visual modality results in a noticeable performance drop across all methods, particularly for FCF (HR@50 decreased by 71.67%) and FedRAP (HR@50 decreased by 30.78%).
-This highlights the critical role of visual information in RecSys, especially for FCF.
-Visual features provide rich, intuitive image-based information that helps models better capture user preferences regarding the appearance of items.
-Without this visual context, models struggle to accurately infer user preferences, leading to a significant decline in recommendation performance.
-Secondly, removing the ID feature also causes a substantial decline in performance.
-The ID feature is essential for personalized recommendations, as it captures users’ historical interactions and reflects their past behaviors and interests.
-When the ID feature is removed, the model loses its primary mechanism for personalization, resulting in decreased accuracy and personalization of recommendations.
-This impact is especially pronounced in methods like PFedRec (HR@50 decreased by 82.82%) and FedRAP (HR@50 decreased by 51.92%), which heavily rely on ID features.
-In contrast, removing the text modality has a relatively smaller impact, though it still affects some methods.
-For instance, in FedAvg and PFedRec, removing the text modality led to HR@50 decreases of 27.63% and 43.21%, respectively.
-This indicates that text information plays an important role in capturing semantic relationships between user behaviors and item descriptions, helping the model better understand user needs and item characteristics, thereby enhancing recommendation performance.
-In summary, all methods experienced performance degradation to varying degrees when any modality was removed.
-Considering Fig.
-2, visual, text, and ID features are often complementary.
-Removing any one of these disrupts the balance, making it difficult for the model to fully represent user preferences and item characteristics.
-Additionally, different methods depend on different modalities to varying extents.
-For instance, FCF and PFedRec are more dependent on visual information, whereas FedAvg is less reliant on ID features.
-These differences in dependency lead to varying levels of performance decline when specific modalities are removed.
-Further experimental results are provided in the appendix.
-Figure 3. Performance comparison of different feature fusion strategies in various FedRec methods combined with FedMR on the KU dataset.
-5.6.3. Ablation Study on Fusion Strategy
+We assess the contribution of visual, textual, and ID feature types by evaluating FedMR-enhanced model performance upon removing each source individually, referencing the updated results in Table 3.
+The findings reveal a clear interdependence among these modalities, confirming their complementary roles.
+ID features remain foundational for personalization, and their removal consistently causes the most significant performance degradation across baseline methods, often proving the most critical factor, particularly impacting models with strong personalization components like FedRAP.
+Visual features are also highly influential, with their absence substantially hindering performance, especially for visually-reliant architectures such as FCF.
+Textual features provide valuable semantic context, and removing them leads to notable accuracy decreases for various models, confirming their complementary contribution beyond visual and ID signals.
+Crucially, we can see the degree of reliance on each modality differs markedly across the underlying FedRec architectures.
+This differential dependency and demonstrated complementarity across modalities necessitate adaptive fusion mechanisms.
+FedMR’s MFFM addresses this requirement, ensuring robust performance by effectively utilizing all available information sources.
+(a) User Distribution
+(b) Modality Preferences
+(c) Heterogeneity Distribution
+(d) Performance Trend
+Figure 3.
+Analysis of user characteristics, heterogeneity, and performance for FedRAP w/ FedMR on the KU dataset.
+(a) Distribution of users across interaction counts; (b) Variation in user preferences across modalities relative to the global average; (c) Distribution of user heterogeneity, which deviation from global preference, across interaction groups; (d) Recommendation performance trend on Recall@50 and NDCG@50 by user interaction group.
+5.6.3. Ablation Study on Client Heterogeneity
+We analyze client heterogeneity and its performance implications using the FedRAP enhanced by FedMR on the KU dataset, with results consolidated in Fig.
+3.
 Fig.
-3 presents the HR@50 and NDCG@50 evaluation results for five used FedRec methods using different feature fusion strategies with FedMR on the KU dataset.
-The results indicate that different fusion strategies significantly impact performance.
-Notably, the Mix strategy, which dynamically combines multiple fusion methods based on user interaction history, outperforms others, while the Sum strategy shows poor performance.
-This trend is consistent across all metrics.
-The Sum strategy simply adds different modality features without considering their complex interactions and inter-relationships.
-Since each modality’s importance varies, simple addition can lead to information loss or imbalanced weighting, making it difficult to capture relationships effectively.
-Additionally, RecSys often face a long-tail distribution, where a few popular items dominate user interactions (Park and Tuzhilin, 2008; Yin et al., 2012).
-The simplicity of the Sum strategy may lead to overfitting on popular items, resulting in poor performance on less interacted items and affecting overall model performance.
-In contrast, the MLP strategy uses fully connected networks to learn nonlinear modality combinations, which better captures complex interactions between item modalities and ID features.
-However, its effectiveness is limited by network depth and the number of parameters.
-On the other hand, the Gate strategy learns adaptive weights to dynamically filter the contribution of each modality, optimizing the importance of different features.
-The Mix strategy integrates all fusion methods and adjusts their weights dynamically based on user interaction history, allowing the model to capture more fine-grained personalized item representations.
-This approach enables more accurate matching of user preferences and effectively captures complex relationships between multimodal features, leading to superior performance.
-This study highlights the importance of personalized item embeddings in FedRec.
-In FedMR, each client dynamically adjusts fusion strategy weights through the Mixing Feature Fusion Module, based on user interaction history, to generate personalized item embeddings, enhancing personalization and accuracy while maintaining user privacy, thus improving overall system performance.
+3(a) shows that the dataset comprises users with varying activity levels.
+A key finding is the significant inter-user heterogeneity in modality preferences, as shown in Fig.
+3(b).
+While preferences related to ID features exhibit relatively low variance, user preferences for text and especially vision features vary considerably more, highlighting the diverse ways users perceive and value different types of multimodal information, posing a challenge for standard fusion approaches.
+Furthermore, as we can see in Fig.
+3(c), analyzing the overall heterogeneity distribution reveals that its variance tends to increase with user activity.
+More active users, therefore, not only deviate from the average preference but also exhibit greater diversity amongst themselves compared to less active users, suggesting that experienced users often develop more distinct and varied multimodal preference patterns, consistent with prior work (Liu et al., 2019; Guo et al., 2024).
+This observed trend of increasing heterogeneity variance among more active users directly correlates with the recommendation performance shown in Fig.
+3(d).
+Recall@50 and NDCG@50 tend to decrease for user groups with higher interaction counts.
+The heightened preference diversity within these active groups makes accurate Top-K prediction intrinsically more difficult.
+Capturing the varied, potentially niche multimodal preferences of highly heterogeneous users presents a significant challenge, contributing to the observed degradation in average performance metrics for these groups.
+This challenge underscores the limitations of non-adaptive fusion methods and highlights the core contribution of FedMR.
+Our framework introduces the MFFM, an innovative client-side mechanism specifically designed for dynamic, personalized adaptation of multimodal representations.
+By allowing each client to tailor the fusion process based on individual interaction history, MFFM directly addresses the crucial challenge of modality preference heterogeneity observed empirically, offering a vital step towards effective personalization in diverse federated multimodal environments.
 Table 4. Performance comparison of FedRec methods combined with FedMR, with and without Gaussian noise (mean=0, variance=0.1) on KU dataset for privacy protection.
@@
 NDCG@50 0.1305 0.1236 5.29% $\downarrow$
-5.6.4. Privacy Protection
-Table 4 presents the experimental results of the used FedRec methods combined with FedMR, comparing the performance with and without adding Gaussian noise (mean=0, variance=0.1) to the gradients of the global components uploaded to the server, to protect user privacy.
-The experiments are conducted on the KU dataset during feature fusion.
-The results indicate that all methods experienced some performance degradation after adding noise, particularly on the NDCG metric, which is more sensitive to ranking accuracy.
-This is because noise disturbs the gradient precision, which affects the model’s ability to update effectively.
-Among the methods, FedNCF is the most sensitive to noise, with HR@50 dropping by 14.33% and NDCG@50 by 31.03%.
-As a complex neural network-based model, FedNCF depends on numerous parameters to capture nonlinear features.
-The introduction of noise disrupts these feature interactions, resulting in a significant decline in performance.
-In contrast, FedAvg and FedRAP show greater robustness to noise, with only minimal performance degradation.
-This suggests that these methods are more effective at maintaining recommendation accuracy while ensuring privacy protection.
-In conclusion, although adding noise helps protect user privacy, its impact varies based on model complexity and feature interaction mechanisms.
-Simpler models like FedAvg and adaptive models like FedRAP are more resilient to noise, while complex neural network models like FedNCF experience more significant performance drops due to the noise interference.
+5.6.4. Ablation Study on Privacy Preservation
+A key consideration in FedRec is compatibility with privacy-enhancing technologies.
+We assess FedMR’s robustness by evaluating the performance impact of adding Gaussian noise to client gradients before server aggregation, a common technique related to differential privacy.
+Table 4 presents these results on the KU dataset.
+As anticipated, introducing noise creates a privacy-utility trade-off, generally leading to decreased recommendation accuracy across all tested FedMR-enhanced models.
+However, the extent of this performance degradation differs significantly depending on the base FedRec architecture.
+Models with simpler structures or strong adaptive components, specifically FedAvg, FCF, and FedRAP, demonstrate considerable resilience, experiencing relatively minor performance reductions.
+Conversely, methods employing more complex neural networks, such as FedNCF, exhibit higher sensitivity to the injected noise, resulting in a more substantial drop in accuracy.
+This differential impact suggests that noise may interfere more significantly with the intricate parameter relationships learned by complex models.
+Nevertheless, the performance reduction remains moderate for most configurations, indicating that FedMR effectively integrates with various baseline methods and maintains reasonable utility under standard noise-based privacy mechanisms, confirming its practical applicability for privacy-conscious recommendation tasks in federated environments.
 6. Conclusion
-This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client.
-This module dynamically adjusts fusion weights based on user interaction history, generating personalized item representations that capture fine-grained user preferences, thereby enhancing personalized recommendations.
-It allows flexible modification of fusion strategies and supports different modalities of item data as needed.
-Experimental results on four multimodal recommendation datasets show that FedMR performs well and is compatible with existing ID-based methods without requiring changes to the underlying framework.
-This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.
+Motivated by privacy concerns and regulations that necessitate alternatives to data-intensive centralized multimodal recommendation, this paper introduced FedMR, a novel framework addressing the critical challenge of effectively integrating multimodal data within privacy-preserving Federated Recommendation Systems.
+FedMR’s core innovation lies in the client-side Mixing Feature Fusion Module, which facilitates dynamic, personalized adaptation by fusing server-provided multimodal representations with ID-based signals according to individual user interaction histories, explicitly modeling and addressing user heterogeneity in multimodal preference fusion directly on the client.
+This adaptive mechanism effectively handles diverse user preferences across modalities.
+Extensive experiments demonstrate that FedMR seamlessly enhances various ID-based FedRecs, substantially boosting their performance and significantly narrowing the gap towards centralized multimodal systems, occasionally even surpassing them.
+By enabling adaptive multimodal fusion under federated constraints, FedMR offers a practical and robust solution, advancing the state-of-the-art towards effective, personalized, and privacy-preserving multimodal recommendation in federated environments.
 Appendix A Algorithm Optimization
@@
 This ensures that $\gamma_{j}^{(u)}$ is updated in an efficient and mathematically sound manner within the FL framework.
+Appendix B Theoretical Analysis
+B.1. Convergence Analysis
+Proof of Lemma 1 (Local Update Bound).
+Using the local update rule:
+(16) $\Theta_{l,u}^{t,k}=\Theta_{l,u}^{t,k-1}-\eta\,g_{l,u}(\Theta_{l,u}^{t,k-1};\xi),$
+and Assumption 2, we have:
+(17) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,k}-\Theta_{l,u}^{t}\Bigr{\|}^{2}&=\mathbb{E}\left\|\sum_{i=1}^{k}\eta\,g_{l,u}(\Theta_{l,u}^{t,i-1};\xi)\right\|^{2}\\[4.0pt] &\leq k\,\eta^{2}\sum_{i=1}^{k}\mathbb{E}\left\|g_{l,u}(\Theta_{l,u}^{t,i-1};\xi)\right\|^{2}\quad\text{(by Jensen's inequality)}\\[4.0pt] &\leq k^{2}\,\eta^{2}(G^{2}+\sigma_{l}^{2})\quad\text{(by Assumption 2).}\end{split}$
+Jensen’s inequality is used because the square of a sum under an expectation can be bounded by $k$ times the sum of squares.
+For $\eta=\tfrac{1}{L\sqrt{A}}$ and $k=A$, it follows that:
+(18) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}&\leq A^{2}\,\eta^{2}(G^{2}+\sigma_{l}^{2})=\frac{A(G^{2}+\sigma_{l}^{2})}{L^{2}}.\end{split}$
+Defining $C_{1}=A$, we can equivalently express the bound as the followings:
+(19) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}\leq C_{1}\,A\,\eta^{2}(G^{2}+\sigma_{l}^{2}).\end{split}$
+Proof of Lemma 2 (Gradient Bias Bound).
+The gradient bias is given by
+(20) $\begin{split}\|B_{t}\|&=\Bigl{\|}\mathbb{E}_{u}\Bigl{[}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}\Bigr{]}-\nabla_{\Theta_{g}}F(\Theta_{g}^{t})\Bigr{\|}\\[4.0pt] &=\Bigl{\|}\mathbb{E}_{u}\Bigl{[}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}-\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t}\bigr{)}\Bigr{]}\Bigr{\|}\\[4.0pt] &\leq\mathbb{E}_{u}\Bigl{\|}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}-\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t}\bigr{)}\Bigr{\|}\\[4.0pt] &\leq L\,\mathbb{E}_{u}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|},\end{split}$
+where the last inequality follows from Assumption 1. Squaring both sides and applying the result from Lemma 1 yields:
+(21) $\begin{split}\|B_{t}\|^{2}\leq L^{2}\,\mathbb{E}_{u}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}\leq L^{2}\,C_{1}\,A\,\eta^{2}(G^{2}+\sigma_{l}^{2}).\end{split}$
+Lemma 0 (Aggregated Gradient Second Moment Bound).
+Under Assumptions 2, 3, and Lemma 2, assuming uniform client sampling on $n_{s}$ clients, we can derive the aggregated gradient second moment bound as followings:
+(22) $\mathbb{E}[\|\bar{g}_{g}^{t}\|^{2}]\leq 2\mathbb{E}[\|\nabla F(\Theta_{g}^{t})\|^{2}]+2\mathbb{E}[\|B_{t}\|^{2}]+\frac{2(\sigma^{2}+G^{2})}{n_{s}}$
+Proof of Lemma 1 (Aggregated Gradient Second Moment).
+When Assumptions 2 and 3 hold, we can use the variance decomposition to derive the followings:
+(23) $\begin{split}\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}&=\Bigl{\|}\mathbb{E}\bar{g}_{g}^{t}\Bigr{\|}^{2}+\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}-\mathbb{E}\bar{g}_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &=\Bigl{\|}\nabla F(\Theta_{g}^{t})+B_{t}\Bigr{\|}^{2}+\operatorname{Var}\left(\bar{g}_{g}^{t}\right)\\[3.0pt] &\leq 2\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+2\|B_{t}\|^{2}+\frac{1}{n_{s}}\,\mathbb{E}_{u}\Bigl{\|}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}\Bigr{\|}^{2}\\[3.0pt] &\leq 2\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+2\|B_{t}\|^{2}+\frac{G^{2}+\sigma^{2}}{n_{s}}.\end{split}$
+The last step is based on Assumptions 2 and 3. ∎
+Proof of Theorem 1 (Convergence Rate).
+Starting from the $L$-smoothness of $F$, which is in Assumpution 1, we have:
+(24) $\begin{split}F(\Theta_{g}^{t+1})&\leq F(\Theta_{g}^{t})+\bigl{\langle}\nabla F(\Theta_{g}^{t}),\Theta_{g}^{t+1}-\Theta_{g}^{t}\bigr{\rangle}+\frac{L}{2}\Bigl{\|}\Theta_{g}^{t+1}-\Theta_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &=F(\Theta_{g}^{t})-\eta\,\bigl{\langle}\nabla F(\Theta_{g}^{t}),\bar{g}_{g}^{t}\bigr{\rangle}+\frac{L\eta^{2}}{2}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}.\end{split}$
+Taking expectations and using Lemmas 2 and 3, we obtain:
+(25) $\begin{split}\mathbb{E}F(\Theta_{g}^{t+1})&\leq\mathbb{E}F(\Theta_{g}^{t})-\eta\,\mathbb{E}\Bigl{\langle}\nabla F(\Theta_{g}^{t}),\nabla F(\Theta_{g}^{t})+B_{t}\Bigr{\rangle}+\frac{L\eta^{2}}{2}\,\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &\leq\mathbb{E}F(\Theta_{g}^{t})-\frac{\eta}{2}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+\frac{\eta}{2}\,\mathbb{E}\|B_{t}\|^{2}\\[3.0pt] &\quad+L\eta^{2}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+L\eta^{2}\,\mathbb{E}\|B_{t}\|^{2}+\frac{L\eta^{2}(G^{2}+\sigma^{2})}{2n_{s}}.\end{split}$
+Here we see the interplay of the gradient norm $\|\nabla F(\Theta_{g}^{t})\|$ and the bias term $B_{t}$.
+We control these via the learning rate $\eta$.
+Choosing $\eta\leq\frac{1}{4L}$ ensures certain terms are upper-bounded by $\frac{1}{2}\|\nabla F(\Theta_{g}^{t})\|^{2}$.
+For $\eta\leq\frac{1}{4L}$, collecting like terms yields:
+(26) $\begin{split}\frac{\eta}{4}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}&\leq\mathbb{E}\Bigl{[}F(\Theta_{g}^{t})-F(\Theta_{g}^{t+1})\Bigr{]}+\frac{5L^{3}C_{1}A\,\eta^{3}(G^{2}+\sigma_{l}^{2})}{2}\\ &\quad+\frac{L\eta^{2}(G^{2}+\sigma^{2})}{2n_{s}}.\end{split}$
+Summing over $t=0$ to $T-1$ and setting $\eta=\frac{1}{L\sqrt{T}}$, we obtain
+(27) $\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}&\leq\frac{4L\,(F_{0}-F^{*})}{\sqrt{T}}+\frac{10C_{1}AG^{2}}{T}+\frac{2(G^{2}+\sigma^{2})}{\sqrt{T}\,n_{s}}.\end{split}$
+The additional term $\mathcal{O}\bigl{(}G^{2}/T\bigr{)}$ is dominated by $\mathcal{O}\bigl{(}G^{2}/\sqrt{T}\bigr{)}$, which concludes the proof. ∎
 Table 5. Changes of parameter numbers in FedRec with and without FedMR on the KU Dataset, Compared to Centralized Recommendation Systems.
@@
 FedMR 742340 1885124 1320900 2757572
-Appendix B Parameter Analysis
+Appendix C Parameter Analysis
 Table 5 illustrates the change in the number of learnable parameters on a single client in the KU dataset when FedMR is integrated into FedRec or not, compared to centralized recommendation systems.
@@
 This is because FedMR incorporates multimodal information (e.g., visual and textual features) alongside ID embeddings, increasing the model’s complexity and consequently requiring more parameters to learn.
-Despite the increase in the number of learnable parameters, Table 2 demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset.
+Despite the increase in the number of learnable parameters, Table 2 of the main paper demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset.
 By combining multimodal features with ID embeddings, FedMR better leverages users’ interaction histories to capture the fine-grained personality of users.
@@
 Thus, the increased learnable parameter size is not only a necessary cost for improved recommendation performance but also a vital measure for ensuring privacy.
-Appendix C More Experiments
+Appendix D More Experiments
+D.1. More Experiment Details
+All baseline methods are implemented in PyTorch (Paszke et al., 2019), and their official implementations are publicly available.
+Experiments are conducted on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, three Quadro RTX 6000 GPUs, and 395 GB of RAM.
+Figure 4. Training loss progression over epochs for baseline methods enhanced with FedMR on the KU dataset, illustrating convergence behavior.
+D.1.1. Convergence Analysis
+To assess the training stability and convergence properties of the baseline methods when enhanced with FedMR, we plot their training loss progression over 50 epochs on the KU dataset in Figure 4.
+The loss curves demonstrate that all FedMR-enhanced models exhibit stable convergence behavior.
+The training loss decreases consistently across epochs for all methods, with a sharp reduction typically observed in the initial epochs followed by more gradual convergence towards a stable value.
+This indicates that the FedMR framework integrates well with diverse base FedRec architectures, allowing for effective optimization and stable learning dynamics.
+While the final converged loss values vary slightly across methods, all models successfully minimize the training objective, confirming the robustness of the training process when incorporating FedMR’s multimodal fusion capabilities.
+Figure 5.
+Sensitivity analysis of 4 FedMR-enhanced baselines (FCF, FedAvg, FedNCF, PFedRec) on the KU dataset.
+Performance is plotted against varying learning rates (lr)and L2 regularization (l2_reg) strengths.
+Figure 6.
+Hyperparameter tuning for FedMR-enhanced FedRAP on the KU dataset.
+Heatmaps visualize HR@50 and NDCG@50 across personalization parameters $\alpha$ and $\beta$ for different learning rates (lr) and L2 regularization (l2_reg).
+Performance intensity is indicated by color.
+D.1.2. Hyperparameter Analysis
+We analyze the sensitivity of FedMR-enhanced models to key hyperparameters, specifically the learning rate and L2 regularization strength, using the KU dataset.
+Fig.
+5 illustrates this for FedMR applied to FCF, FedAvg, FedNCF, and PFedRec.
+Across these models, the learning rate exhibits a significant impact, with performance generally peaking at intermediate values and decreasing notably at lower or higher rates.
+This highlights the necessity of selecting an appropriate learning rate for effective convergence.
+The influence of L2 regularization appears moderate in comparison, though still relevant; no single $l2\_reg$ value is universally optimal, with the best performance achieved at different strengths depending on the base model and learning rate, necessitating model-specific tuning.
+The tuning process for the FedMR-enhanced FedRAP model, which includes its specific personalization parameters $\alpha$ and $\beta$, is detailed in Fig.
+6.
+The heatmaps clearly indicate high sensitivity to the choices of $\alpha$ and $\beta$, with distinct optimal regions emerging.
+Furthermore, these optimal regions shift based on the learning rate, signifying a strong interaction between $lr$, $\alpha$, and $\beta$.
+This underscores the need for careful co-tuning, typically via grid search, to identify the best configuration for FedMR-enhanced FedRAP.
+Overall, while the models demonstrate reasonable robustness, achieving optimal results requires careful hyperparameter tuning, particularly for the learning rate and any model-specific parameters.
+The best hyperparameters identified through validation during this analysis informed the settings used for the main experimental results.
+Figure 7. Visualization of multimodal features and ID features mapped into a shared representation space.
+D.1.3. Ablation Study on Modality Distributions
+To analyze the interaction between multimodal features and ID features, we use FedRAP as the backbone for FedMR.
+Fig.
+7 visualizes the distribution of these features on the KU dataset.
+The first subfigure shows that globally shared ID features and multimodal features occupy distinct subregions in the shared representation space, a phenomenon known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a).
+This separation enhances the robustness of multimodal data processing by reducing noise interference.
+The subsequent subfigures illustrate how personalized ID features influence this distribution.
+Incorporating personalized ID embeddings (e.g., for Users 00 and $42$) causes the ID features to expand toward the regions of visual and textual features, effectively shrinking the modality gap.
+This adjustment facilitates multimodal fusion and improves the accuracy of personalized recommendations.
+By comparing users, it is evident that the model dynamically adapts feature distributions based on user-specific behaviors and preferences, better addressing individual needs.
+This demonstrates the critical role of personalized ID features in federated recommendation systems, enabling effective personalization while preserving user privacy.
 Figure 8. Performance comparsions of different FedRec methods with FedMR under varying client sampling ratios on HR@50 and NDCG@50.
-C.1. Ablation Study on Client Sampling
+D.2. Ablation Study on Client Sampling
 To investigate the impact of introducing FedMR for multimodal data fusion on different FedRec methods under varying client sampling ratios, we conduct experiments on the five FedRec methods using the KU dataset.
@@
 This is especially true in multimodal scenarios, where a lack of sufficient modality information weakens the fusion effect, reducing recommendation accuracy.
-Appendix D Limitations and Future Work
-While FedMR has shown promising improvements in federated multimodal recommendation systems, several limitations remain:
-  1. (1)
-    Data Heterogeneity: Considering the federated settings, the distribution of user data varies significantly.
-Although FedMR mitigates this issue by integrating multimodal information, its performance is still limited when handling highly heterogeneous data.
-Future work could focus on introducing more robust personalization models or adaptive aggregation strategies to better address the differences between users.
-  2. (2)
-    Communication Overhead: Although FedMR reduces some communication costs by encoding multimodal data, processing multimodal features still results in high computational and communication costs.
-Future research could explore optimizing communication protocols or using model compression techniques to further minimize these costs.
-  3. (3)
-    Limited Flexibility in Multimodal Fusion: While FedMR supports dynamic adjustments to its fusion strategies, its current approach may lack flexibility when dealing with complex user behaviors.
-Future work could explore more adaptive and layered fusion methods to better capture the diverse preferences of users.
-In conclusion, future developments could enhance the overall performance of federated recommendation systems by improving the handling of data heterogeneity, reducing communication costs, and optimizing multimodal fusion strategies.
+Table 6. Performance comparison on the KU dataset. The table reports HR@50 and NDCG@50 for several FedRecs. Percentages indicate the performance drop relative to the full-feature baseline.
+Method Metric FedMR w/o V w/ rnd V w/o C w/ rnd C w/o D w/ rnd D
+FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3922 (23.80% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.3676 (28.58% $\downarrow$) 0.5094 (1.03% $\downarrow$) 0.4520 (12.18% $\downarrow$)
+NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1303 (12.61% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1240 (16.83% $\downarrow$) 0.1407 (5.66% $\downarrow$) 0.1375 (7.78% $\downarrow$)
+FCF HR@50 0.5539 0.2843 (48.67% $\downarrow$) 0.2892 (47.79% $\downarrow$) 0.4216 (23.89% $\downarrow$) 0.2451 (55.75% $\downarrow$) 0.2353 (57.52% $\downarrow$) 0.2206 (60.17% $\downarrow$)
+NDCG@50 0.1597 0.0748 (53.16% $\downarrow$) 0.0783 (50.97% $\downarrow$) 0.0965 (39.57% $\downarrow$) 0.0629 (60.61% $\downarrow$) 0.0578 (63.81% $\downarrow$) 0.0538 (66.31% $\downarrow$)
+FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1667 (14.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1618 (17.49% $\downarrow$) 0.1602 (18.31% $\downarrow$) 0.1569 (19.99% $\downarrow$)
+NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0489 (19.70% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0461 (24.30% $\downarrow$) 0.0538 (11.66% $\downarrow$) 0.0458 (24.79% $\downarrow$)
+PFedRec HR@50 0.2157 0.1990 (7.74% $\downarrow$) 0.1647 (23.64% $\downarrow$) 0.1941 (10.01% $\downarrow$) 0.1696 (21.37% $\downarrow$) 0.1794 (16.83% $\downarrow$) 0.1647 (23.64% $\downarrow$)
+NDCG@50 0.0616 0.0469 (23.86% $\downarrow$) 0.0468 (24.03% $\downarrow$) 0.0537 (12.82% $\downarrow$) 0.0445 (27.76% $\downarrow$) 0.0473 (23.21% $\downarrow$) 0.0484 (21.43% $\downarrow$)
+FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2206 (42.31% $\downarrow$) 0.3235 (15.40% $\downarrow$) 0.2304 (39.75% $\downarrow$) 0.1863 (51.29% $\downarrow$) 0.1324 (65.38% $\downarrow$)
+NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0634 (51.42% $\downarrow$) 0.0714 (45.29% $\downarrow$) 0.0564 (56.78% $\downarrow$) 0.0490 (62.43% $\downarrow$) 0.0338 (74.10% $\downarrow$)
+D.2.1. Ablation Study on Modality Contribution
+We present the full modality contribution analysis, including results for noise replacement ($w/rnd$), in Table 6 owing to page limits.
+We assess the individual contributions of visual ($\mathbf{V}$), textual ($\mathbf{C}$), and ID ($\mathbf{D}$) feature types within FedMR by evaluating performance upon either removing the feature type entirely ($w/o$) or replacing it with random noise ($w/rnd$).
+Table 6 presents these results on the KU dataset, revealing the distinct and complementary roles of each modality.
+The comparison between feature removal and noise replacement elucidates the nature of each modality’s contribution.
+ID features consistently prove indispensable for personalization.
+Their removal ($w/o~{}D$) incurs substantial performance degradation across all methods.
+Notably, replacing ID features with random noise ($w/rnd~{}D$) often leads to even more severe performance drops, particularly for models with strong personalization components like FCF and FedRAP.
+This trend underscores that models heavily rely not just on the presence of an ID signal, but on the specific collaborative and structural information learned within the ID embeddings.
+Visual features also demonstrate significant influence.
+Removing them ($w/o~{}V$) markedly hinders performance, especially for architectures known to leverage visual cues, such as FCF and FedRAP.
+Comparing this to noise replacement ($w/rnd~{}V$), we observe that substituting visual features with noise frequently results in equal or greater performance degradation than simple removal.
+This finding suggests that visual features provide meaningful, non-random signals that benefit recommendation, and replacing them with noise is often more detrimental than their absence, particularly for visually sensitive models.
+Textual features provide valuable complementary semantic context.
+While their removal ($w/o~{}C$) impacts performance to a lesser extent than ID or visual features for some models, the effect is still notable across various methods.
+Replacing textual features with noise ($w/rnd~{}C$) generally causes a more significant performance decrease than simple removal, confirming that the semantic content captured by text offers a non-trivial contribution beyond just acting as a placeholder.
+Collectively, these results highlight a clear interdependence among ID, visual, and textual information sources.
+The consistent performance drop upon removal or noise replacement confirms that each modality contributes uniquely.
+The often more detrimental effect of noise replacement compared to feature removal strongly indicates that the learned multimodal features contain meaningful signals crucial for performance.
+Furthermore, the degree of reliance on each modality, and the differential impact of removal versus noise replacement, varies markedly across the underlying FedRec architectures.
+This complex interplay and differential dependency motivate the need for adaptive fusion mechanisms.
+FedMR’s MFFM addresses this by dynamically balancing the contributions from these distinct and valuable information sources based on user context, enabling robust performance through effective information leverage.
+Appendix E More Related Work
+E.1. Federated Recommendation Systems
+Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020).
+In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024b), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d).
+However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have.
+AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec.
+AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features.
+In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy.
+Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks.
+Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features.
+FedDAE (Li et al., 2024c) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network.
+While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
+E.2. Multimodal Recommendation Systems
+Unlike traditional Recommendation Systems (RecSys) that rely primarily on user-item interaction data (such as clicks or ratings) to capture user behavior (Yu et al., 2023), Multimodal Recommendation systems (MMRec) aim to improve the accuracy of recommendations and enhance user experience by utilizing multiple modalities of data , such as text, images, audio, and video (Zhou et al., 2023b).
+As unique identifiers for items, item IDs not only contain semantic information about the items but also capture the characteristics of neighboring items (Yuan et al., 2023b; Liu et al., 2023).
+Therefore, in MMRec, item IDs are often used as auxiliary information and combined with multimodal data to enhance item representations(Wang et al., 2024; Zhang et al., 2024e).
+This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving the quality of recommendations.
+A typical MMRec training process generally involves feature extraction and modality fusion (He and McAuley, 2016; Wei et al., 2019; Wang et al., 2021; Tao et al., 2022).
+Feature extraction (Guo et al., 2019) aims to obtain representative features from multiple modalities to better capture the attributes and preferences of users and items.
+Modality fusion (Gao et al., 2020) focuses on effectively integrating data from different modalities to enhance the overall performance of the MMRec and the relevance of its results.
+The recommendation models (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023) combine multimodal data with user-item interaction data to predict user preferences and provide personalized recommendations.
+E.3. Foundation Models
+Foundation Models (FM) are often pre-trained on large-scale unlabeled datasets for general purposes, that can be adapted to a variety of downstream tasks through appropriate fine-tuning, such as task-specific training (Bommasani et al., 2021).
+This makes them highly generalizable and adaptable across different applications.
+They are then adapted to specific recommendation tasks through methods such as representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), model fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), and prompt tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024).
+Some studies have explored the use of FM to enhance the content-awareness of MMRec, improving content understanding and generation by leveraging alignment and generation techniques across multiple modalities.
+CLIP (Radford et al., 2021a; Wu et al., 2024) uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain multimodal representations, capturing the matching relationships between users and items.
+M6-Rec (Cui et al., 2022) builds a recommendation system based on large-scale multimodal pre-trained models specifically designed for the e-commerce domain.
+P5 (Geng et al., 2022) transforms the recommendation task into a natural language processing task, generating personalized recommendations through prompts.
+VIP5 (Geng et al., 2023) leverages pre-trained multimodal models to capture complementary information from visual and textual data, providing more personalized and relevant recommendations to users.
+FedPA (Zhang et al., 2024c) and TransFR (Zhang et al., 2024b) integrate FM with FedRec.
+FedPA introduces a personalized adapter to fine-tune the FM, allowing each client to learn individual user preferences in a lightweight manner.
+In contrast, TransFR leverages FMs to generate universal textual representations, replacing traditional ID-based item representations and enhancing the system’s adaptability to new domains and its ability to handle cold-start issues.
+Appendix F Limitations and Future Work
+While FedMR demonstrates promising results for privacy-preserving multimodal recommendation, key limitations suggest avenues for future research.
+The specific subfield of FedMMRec is in its early stages: while FedMR advances this area alongside recent works like AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a), establishing direct experimental comparisons to precisely gauge relative performance remains a key challenge for future work as the field evolves and common benchmarks emerge.
+Additionally, FedMR primarily utilizes visual and textual data via CLIP, future research should explore incorporating a broader range of relevant modalities, such as audio or temporal information from videos.
+Further investigations could also focus on enhancing the client-side MFFM personalization mechanism or integrating FedMR with advanced FL techniques to better address communication constraints or statistical heterogeneity.